### v1

# Rubrics as Rewards: Reinforcement Learning Beyond Verifiable Domains

Abstract
Extending Reinforcement Learning with Verifiable Rewards (RLVR) to real-world tasks often requires balancing objective and subjective evaluation criteria. However, many such tasks lack a single, unambiguous ground truth—making it difficult to define reliable reward signals for post-training language models. While traditional preference-based methods offer a workaround, they rely on opaque reward functions that are difficult to interpret and prone to spurious correlations. We introduce Rubrics as Rewards (RaR), a framework that uses structured, checklist-style rubrics as interpretable reward signals for on-policy training with GRPO. Our best RaR method yields up to a 28% relative improvement on HealthBench-1k compared to simple Likert-based approaches, while matching or surpassing the performance of reward signals derived from expert-written references. By treating rubrics as structured reward signals, we show that RaR enables smaller-scale judge models to better align with human preferences and sustain robust performance across model scales.
Rubrics as Rewards:
Reinforcement Learning Beyond Verifiable Domains
Anisha Gunjal   Anthony Wangthanks: Work done while at Scale AI.   Elaine Lau Vaskar NathBing LiuSean Hendryx Scale AI anisha.gunjal@scale.com
1 Introduction
Reinforcement Learning with Verifiable Rewards (RLVR) has enabled large language models to elicit complex reasoning on tasks with verifiable outcomes, most prominently in mathematics and coding, where a reward model can be replaced with a scoring function and final answers are easily verifiable or test cases provide automated mechanisms for evaluation Lambert et al. (2024); Guo et al. (2025a); Cui et al. (2025). However, many real-world tasks lack such explicit verifiable answers, leaving models without direct reward feedback. In practice, researchers often turn to RLHF via preference ranking—collecting human judgments over pairs or lists of model outputs to fill this gap. While preference-based reward models can bootstrap performance, they tend to overfit superficial artifacts (e.g. response length, formatting quirks, annotator biases) Singhal et al. (2023); Wang et al. (2024); Chen et al. (2024b); Ye et al. (2024); Gudibande et al. (2023) and require large volumes of pairwise comparisons Ouyang et al. (2022), making preference-based reward models brittle and costly.
To address this gap, we introduce Rubrics as Rewards (RaR) for on-policy Reinforcement Learning that treats structured criteria or rubrics as the core reward mechanism. We use rubrics as checklists Arora et al. (2025); Sirdeshmukh et al. (2025) composed of independent subgoals allowing for automatable feedback aligned with expert intent. By decomposing “what makes a good response” into tangible, human-interpretable criteria, rubrics offer a middle ground between binary correctness signals and coarse preference rankings.
Figure 1: Overview of Rubrics as Rewards (RaR). (i) Rubric Generation: We synthesize prompt-specific, self-contained rubric criteria using a strong LLM guided by four core design principles, with reference answers serving as proxies for expert supervision. (ii) GRPO Training: These rubrics are used to prompt an LLM judge for reward estimation, which drives policy optimization via the GRPO on-policy learning loop.
Previous works train generative reward models that learn to evaluate reasoning or final outputs with interpretable scores Chen et al. (2025); Whitehouse et al. (2025); Anugraha et al. (2025); Guo et al. (2025b), and some have even used a model’s internal confidence estimates as a proxy for reward Zhao et al. (2025). Concurrently, recent efforts have extended verifiable datasets beyond STEM domains, broadening the applicability of RLVR methods to a wider range of tasks Su et al. (2025b); Ma et al. (2025). However, developing a general-purpose approach for specifying reliable and scalable rewards remains an open challenge, particularly in settings where no single ground truth exists, or where both subjective and objective criteria must be considered to evaluate response quality. The Rubrics as Rewards strategy offers a flexible solution by repurposing structured evaluation criteria into multi-dimensional reward signals. Figure 1 illustrates our approach for generating rubrics and leveraging them as reward signals for on-policy training.
Our key contributions are as follows: (i) We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning framework that uses checklist-style rubrics to supervise multi-criteria tasks, enabling stable training and improved performance in both reasoning and real-world domains. (ii) We apply our rubric generation approach to the medicine and science domains, producing two training datasets—RaR-Medicine-20k and RaR-Science-20k [^1] The datasets will be released soon.. (iii) We show that models trained with RaR match or outperform strong baselines, achieving notable gains in accuracy across diverse domains. (iv) By using rubrics as structured reward signals, RaR enables smaller judge models to achieve better alignment with human preferences and maintains robust performance across different model scales.
2 Rubrics as Rewards
2.1 Problem Formulation
Let $x$ denote an input prompt and $\hat{y}\sim\pi_{\theta}(\cdot\mid x)$ be a sampled response from a model parameterized by $\theta$. In domains without single ground-truth answers or automatic correctness signals, we define a structured reward function using prompt-specific rubric criteria.
Each prompt $x$ is associated with a set of $k$ rubric items $\{(w_{j},c_{j})\}_{j=1}^{k}$, where $w_{j}\in\mathbb{R}$ denotes the weight of criterion $j$, and $c_{j}:(x,\hat{y})\mapsto\{0,1\}$ is a binary correctness function that indicates whether the response $\hat{y}$ satisfies that criterion given the prompt. The final normalized scalar reward is computed as:
$r(x,\hat{y})=\frac{\sum_{j=1}^{k}w_{j}\cdot c_{j}(x,\hat{y})}{\sum_{j=1}^{k}w_{j}}$ (1)
This normalization ensures reward values are consistent across prompts with varying numbers and weights of rubric items.
Explicit Rubric Aggregation.
Each criterion is evaluated independently using a generative reward model, and the reward is computed using Equation 1. While we currently assume binary values for $c_{j}$, this framework supports extension to continuous-valued judgments.
Implicit Rubric Aggregation.
An alternative approach passes all rubric descriptions and weights to a generative reward model, which internally computes a scalar reward:
$r_{\text{implicit}}(x,\hat{y})=f_{\phi}(x,\hat{y},\{(w_{j},d_{j})\}_{j=1}^{k})$ (2)
Here, $f_{\phi}$ denotes an LLM-based judge that takes as input the prompt $x$, the generated response $\hat{y}$, and a set of weighted rubric criteria. Each pair $(w_{j},d_{j})$ consists of a scalar or categorical weights $w_{j}\in\mathbb{R}$ and a rubric criterion $d_{j}$ describing a specific aspect of response quality. This formulation delegates the aggregation of evaluation criteria to the model itself, allowing it to compute a holistic reward score.
In practice, we use an LLM-as-judge with separate prompt templates containing prompt, response, and rubric items for explicit and implicit aggregation. The prompts used in our experiments are detailed in Appendix E.
2.2 Generalization of RLVR with Rubrics as Rewards
Rubric-based reinforcement learning extends the standard RLVR (Reinforcement Learning with Verifiable Rewards) setting by supporting multi-dimensional, prompt-specific evaluation criteria. We formalize this relationship below.
Remark 1 (Rubrics as Rewards subsumes RLVR).
[…]
where $\mathrm{match}(y,\hat{y})\in\{0,1\}$ indicates whether the response satisfies the verifiable correctness condition.
Rubric-based reward functions thus generalize RLVR by enabling multi-dimensional supervision, flexible weighting across criteria, and the incorporation of both objective and subjective aspects of response quality.
This formalization highlights that RLVR can be seen as a restricted instance of rubric-guided RL with a single essential criterion. In contrast, rubric-based rewards further enable structured supervision in settings where correctness is multifaceted and may not be strictly verifiable.
3 Rubric Generation
3.1 Desiderata
We synthesize prompt-specific rubrics guided by a set of design principles. Each rubric defines the criteria for a high-quality response and enables human-interpretable supervision. We use large language models (LLMs) as expert proxies to generate these rubrics while ensuring adherence to the following desiderata:
Grounded in Expert-guidance
Reference answers produced by human experts or stronger LLMs serve as proxies for expert guidance. Rubrics are grounded in these references to capture the key facts, reasoning steps, and conclusions necessary for correctness.
Comprehensive coverage
Rubrics are designed to span multiple quality dimensions, including factual accuracy, logical structure, completeness, style, and common pitfalls. Negative “pitfall” criteria help penalize frequent or high-risk errors. The total number of rubric items per rubric (typically 7–20) is chosen to balance coverage with signal sparsity.
Semantic weighting
Each criterion is labeled with a categorical importance level (e.g., Essential, Important, Optional, Pitfall) that reflects its relative priority in the final reward. This supports interpretable reward aggregation. Future work may explore numeric or learned weighting for greater flexibility.
Self-contained evaluation
Rubrics are written to be independently actionable-each item can be evaluated in isolation by either human annotators or LLM-based judges, without requiring external context or domain knowledge.
3.2 Synthesis
We apply the above framework on real world science and Medicine reasoning domains. For each domain, the prompt (included in Appendix H) instructs the LLM to generate 7–20 rubric items based on the complexity of the input question. Each item is assigned a categorical weight (e.g., Essential Criteria, Important Criteria) to determine its importance to a correct answer. The rubrics are designed to be fully self-contained which means that non-expert readers should be able to evaluate response quality using only the rubric. The evaluation outcome can be directly used as reward functions during reinforcement learning.
We generate synthetic rubrics using OpenAI’s o3-mini and gpt-4o models OpenAI o3-mini (2025); Jaech et al. (2024); Hurst et al. (2024) as described in Section 3. The RaR-Medical-20k and RaR-Science-20k are released for public use. For both domains, we use the reference answers provided in the datasets as a proxy for human expertise while generating synthetic rubrics. These rubric sets are then used to supervise the training of smaller models through either explicit or implicit reward aggregation (Section 2.1).
4 Experiments
Domain $\rightarrow$ Medicine Science
Method/Baseline $\downarrow$ HealthBench-1k GPQA_Diamond
Qwen2.5-7b 0.0818 0.3030 $\pm$ 0.0286
Qwen2.5-7b-Instruct 0.2359 0.3598 $\pm$ 0.0077
Simple-Likert 0.2489 0.3409 $\pm$ 0.0104
Reference-Likert 0.3155 0.3775 $\pm$ 0.0350
Predefined RaR 0.2472 0.3485 $\pm$ 0.0365
RaR-Explicit (o3-mini rubrics) 0.2559 0.3333 $\pm$ 0.0504
RaR-Explicit (GPT-4o rubrics) 0.2979 0.3030 $\pm$ 0.0197
RaR-Implicit (o3-mini rubrics) 0.3107 0.3864 $\pm$ 0.0407
RaR-Implicit (GPT-4o rubrics) 0.3194 0.3662 $\pm$ 0.0191
Table 1: Performance of baseline and RaR (Rubrics as Rewards) methods on medicine and science domains. For HealthBench-1k, we report the overall score; for GPQA_Diamond, we report mean accuracy over 4 runs with standard deviation. All experiments use gpt-4o-mini as the judge model. (RaR) methods are trained using rubrics generated by either o3-mini or GPT-4o. Across both domains, the RaR-Implicit method consistently outperforms Simple-Likert and matches or exceeds the performance of Reference-Likert.
4.1 Datasets
We investigate the utility of rubrics as rewards across two reasoning domains medicine and science.
  • RaR‑Medical‑20k: We curate a set of 20k prompts from medicine‑related subsets of medical‑o1‑reasoning‑SFT Chen et al. (2024a), the natural_reasoning dataset Yuan et al. (2025), the SCP‑116K dataset Lu et al. (2025), and the GeneralThought‑430K dataset General Reasoning (2025).
  • RaR‑Science‑20k: We curate roughly 20k science prompts aligned with the GPQA Diamond benchmark from the natural_reasoning dataset Yuan et al. (2025), the SCP‑116K dataset Lu et al. (2025), and the GeneralThought‑430K dataset General Reasoning (2025).
Refer Tables 7 and 10 in Appendix for the full topic distributions.
4.2 Training Details
We train all models using the GRPO algorithm Shao et al. (2024), with Qwen2.5-7B as the base policy. Training is conducted with a batch size of 96, a learning rate of $5\times 10^{-6}$, and a constant learning rate schedule with 10% linear warmup. Full hyperparameter details are provided in Appendix 11. All models are trained on a single compute node equipped with 8 NVIDIA H100 GPUs.
Our training pipeline consists of the following key components:
Response Generation: For each prompt $q$, we sample $k=16$ responses from the current policy $\pi_{\theta}$, using a context length of 3584 and a sampling temperature of 1.0.
Reward Computation: We use gpt-4o-mini as the judge model to assign rewards $R_{q}$ to the sampled responses. The reward functions are described in Sections 4.3 and 4.4.
Policy Update: The policy is updated using GRPO based on the computed rewards.
4.3 Baselines
We consider compare various baselines and off-the-shelf post-trained models. All rubric-free baselines are with Qwen2.5-7b as the base policy.
For off-the-shelf baselines we evaluate performance on Qwen2.5-7b. We also include the performance of Qwen2.5-7b-Instruct to compare with instruction tuned variant of the base policy.
We describe the post-trained baselines without rubrics below:
SIMPLE-LIKERT:
LLM-judge is asked to output a Likert Score from 1-10 (normalized to 0-1) for each response-prompt pair.
REFERENCE-LIKERT:
The judge model compares the generated response against a high-quality reference answer and outputs a Likert score from 1–10, which is then normalized to the $[0,1]$ range. The reward for each (prompt, response, reference) triplet is defined as:
$R_{\text{ref}}(q,x)=\mathrm{Norm}(\text{LikertScore}(q,x,x^{*}))$
where $x^{*}$ denotes the reference answer, either written by a human or generated by a stronger LLM.
4.4 Rubric-guided Methods
PREDEFINED-RaR:
This method uses a fixed set of generic rubrics (not prompt-specific, as described in Appendix C) and computes the reward via an explicit weighted sum, following Equation 1. All rubric criteria are weighted uniformly.
RaR-EXPLICIT:
This variant also employs explicit aggregation (1) but uses prompt-specific rubrics generated by the method in Section 3. Since the generated rubrics include categorical labels, we assign numerical weights using a simple default scheme: {"Essential": 1.0, "Important": 0.7, "Optional": 0.3, "Pitfall": 0.8}, with higher penalty applied to violated pitfalls.  [^2] Pitfall criteria are phrased in positive form (e.g., “The response avoids misinformation”), so satisfying them contributes positively to the score. If a pitfall is not satisfied, the corresponding reward is reduced or penalized.
RaR-IMPLICIT:
This variant also leverages prompt-specific rubrics; however, instead of aggregating criterion scores explicitly, it prompts the judge model to evaluate the response holistically and assign a single Likert rating (1–10). This rating is then normalized to the $[0,1]$ range:
$R_{\text{implicit}}(q,x)=\mathrm{Norm}(\text{LikertScore}(q,x,R_{q}))$
where $R_{q}$ denotes the list of rubrics associated with prompt $q$.
4.5 Evaluation Setup
Medical Reasoning
We evaluate models trained on curated medical reasoning data using HealthBench-1k Arora et al. (2025), a 1,000-example subset of HealthBench. [^3] We introduce HealthBench-1k as a cost-effective evaluation split for medical reasoning tasks and will release the test set identifiers to support reproducibility.. We generate model responses with greedy decoding (temperature=0) and report the overall_score.
Science Reasoning
We evaluate models trained with curated science set on the GPQA-Diamond subset Rein et al. (2024). Each evaluation is independently repeated 4 times to account for variance, where each run samples one response per prompt using greedy decoding (temperature=0). Answer choices are permuted per example, and model outputs are parsed for boxed answer formats (e.g.,
boxed{A}). If extraction fails, we fall back to a GPT-4o verifier that checks whether the response contains the correct option letter or text (see F). Final accuracy is reported as the mean of the 4 independent runs.
5 Results
We now present the main findings of our study.
Rubrics as Rewards Combine Interpretability with Strong Performance.
Table 1 reports results on the Medicine (HealthBench-1k) and Science (GPQA_Diamond) benchmarks. Our method, RaR-Implicit, consistently outperforms baselines such as Simple-Likert, with the best variant achieving upto 28% relative improvement on HealthBench-1k and 13% on GPQA. It also surpasses both base and instruction-tuned policy models, demonstrating the effectiveness of rubric-guided training for nuanced response evaluation.
Furthermore, RaR-Implicit matches or exceeds the performance of the Reference-Likert baseline. In our setup, synthetic rubrics are generated using reference answers as proxies for expert supervision, meaning their quality is inherently constrained by the completeness and clarity of those references. Despite this, transforming open-ended answers into structured rubric criteria produces reward signals that are both effective and aligned with evaluation objectives.
We also find that implicit aggregation outperforms explicit, weighted summation which indicates that allowing an LLM-based judge to infer how to balance rubric criteria per prompt is more effective than relying on fixed, hand-crafted weights. While explicit aggregation offers more transparency and control, implicit methods better adapt to real-world complexity.
Rubric-based supervision further enables transparent reward specification by breaking response quality into interpretable criteria. Unlike holistic reference comparisons, rubrics reduce stylistic bias and surface-level overfitting. Though reference-based Likert scoring may perform slightly better in some cases, rubrics offer a more scalable and controllable approach—particularly where references are scarce or model interpretability is essential.
Figure 2: LLM Judge Alignment Across Model Scales. Rubric-guided evaluation (orange) consistently improves judge accuracy across model sizes compared to pure Likert-based scoring (blue). Rubrics without reference answers (green) perform better than Likert baseline, though still lag behind full rubric guidance. Smaller scale models benefit the most from rubric structure, narrowing the performance gap with larger judges.
Rubrics enhance alignment with human preferences across model scales
To evaluate the effectiveness of rubric-based evaluations in reflecting human preferences, we generated datasets consisting preferred (chosen) and perturbed (rejected) responses (see Appendix I for details on dataset generation). We use LLM-Judges to assign ratings from 1 to 10 for these responses both with (RaR-IMPLICIT ) and without the guidance of rubrics (SIMPLE-LIKERT). Figure 2 shows how often judges correctly rated the preferred responses higher than the rejected ones across various model scales (from 3b to 72b to o3-mini). The results consistently show that rubric-guided evaluations provide clearer and more accurate signals, as evidenced by the higher accuracy with which preferred responses receive higher ratings when rubrics are involved. This indicates that rubrics improves the alignment with human preferences by offering detailed, context-specific criteria that help judges better discern subtle differences in quality compared to evaluations using only simple Likert-based ratings. Further study on LLM-Judge scaling impacts on policy training are discussed in Appendix C.
Expert guidance is crucial for synthetic rubric generation
Human guidance significantly influences the effectiveness of rubrics in capturing subtle human preferences. Figure 2 highlights performance differences between rubric-based evaluations that include reference answers and those without them. The data shows that rubrics developed with reference answers achieve higher accuracy, emphasizing that human insights integrated during rubric generation enable granular criteria and improved alignment with human preferences.
6 Ablations
Elements of Rubric Design
This ablation study examines how the structure and weighting of synthetic rubrics—generated using o3-mini—affect downstream performance on HealthBench. As shown in Table 2, predefined generic rubrics substantially underperform compared to prompt-specific ones, underscoring the importance of contextualization. Rubrics that include a broader range of criteria—both positive and negative—consistently outperform those limited to essential checks, suggesting that richer evaluation signals lead to better learning. Interestingly, we observe minimal performance differences when including rubric weights or pitfall criteria during training. One possible explanation is that synthetically generating effective pitfall criteria is inherently difficult, as it requires anticipating the most common or critical failure modes of the model, a task that often demands human intuition and domain expertise. As a result, these synthetic negative criteria may lack the specificity or relevance needed to meaningfully penalize undesirable responses.
Ablation Setting Overall Score
Predefined Generic Rubrics 0.2898 $\pm$ 0.00526
Essential-Only Rubrics 0.3562 $\pm$ 0.00976
No Weights 0.3756 $\pm$ 0.00479
No-Pitfall Criteria 0.3837 $\pm$ 0.00420
All Rubrics 0.3723 $\pm$ 0.00005
Table 2: Ablation results on HealthBench-1k (trained on HealthBench-3.5k subset with Qwen2.5-7B base policy). Rubrics generated using o3-mini with access to reference answers. Variations test the impact of rubric structure and weighting on implicit reward aggregation.
Training Method Overall Score
Expert-Answer-SFT 0.204
Simple-Likert 0.239
Reference-Likert 0.317
RaR-Implicit-Synthetic-NoRef 0.320
RaR-Implicit-Synthetic 0.359
RaR-Implicit-Human 0.348
Table 3: Evaluation on HealthBench: Comparison of human- vs. synthetic-generated rubrics (with and without reference answers). RaR methods trained with GRPO significantly outperform Likert-only, Reference-based-likert and SFT baselines. Synthetic rubrics generated without access to reference answers perform notably worse, highlighting the importance of human-grounded guidance. Notably, human-authored rubrics and synthetic rubrics with access to references yield comparable performance.
Impact of Rubric Generation Strategies in Real-World Domains
How does the method of rubric generation affect downstream training in challenging, real-world settings? To investigate this, we evaluate on the HealthBench dataset using a curated subset of 3.5k human-annotated examples, split into train and test sets. Table 3 summarizes the results.
We observe that all rubric-based methods outperform rubric-free baselines, including Reference-Likert. Notably, even the weakest RaR variant significantly surpasses Reference-Likert, underscoring the advantage of structured supervision in subjective, open-ended domains like healthcare. We attribute this to the finer granularity and clarity rubrics provide in assigning rewards-especially when correctness is not binary and answers may vary in tone, completeness, or safety relevance.
Moreover, we find that rubric quality is crucial: synthetic rubrics generated with reference-answer guidance consistently outperform those generated without it. This highlights the importance of incorporating expert signal, whether via human-in-the-loop annotations or high-quality reference completions, for generating effective and aligned rubrics. Purely synthetic rubrics, while scalable, currently fall short in capturing the subtle criteria required for robust training in high-stakes domains.
Impact of LLM Expertise on Rubric Quality
To assess how the capabilities of rubric-generating LLMs affect downstream performance, we generate synthetic rubrics without access to reference answers and use them to train policies on HealthBench. This isolates the effect of LLM quality on reference-free rubric utility. Specifically, we evaluate on the HealthBench-1k subset, using models trained on rubrics generated for the remaining 4k training examples from HealthBench.
As shown in Table 4, larger or more capable LLMs generally produce more effective rubrics, with GPT-4o yielding the best performance among reference-free models. However, all remain below the performance of rubrics generated with reference guidance (e.g., O3-mini with access to reference answers). Additionally, model attributes such as instruction tuning and reasoning capabilities play a key role in the effectiveness of rubric generation.
Rubric Generation Model Overall Score
O3-mini (Rubrics with reference) 0.359
GPT-4o 0.342
GPT-4o-mini 0.327
O3-mini 0.324
Qwen-72B-Instruct 0.327
Qwen-32B-Instruct 0.311
Qwen-7B-Instruct 0.319
Table 4: Policy performance on HealthBench-1k when trained with GRPO using rubrics generated by different LLMs without reference answers. GPT-4o-generated rubrics yield the strongest performance, though they still fall short of rubrics generated with expert (reference-guided) supervision. Smaller aligned models (e.g., GPT-4o-mini, O3-mini) remain competitive with larger open-weight models, underscoring the importance of alignment and reasoning ability in rubric generation.
7 Related Work
RLVR across domains
Reinforcement learning with verifiable rewards (RLVR) is quickly extending its reach well beyond math and code. General-Reasoner trains on a 200 k mixed corpus-physics, finance, policy, and more-and reports a ten-point jump on MMLU-Pro after GRPO fine-tuning Ma et al. (2025). A follow-up study extends RLVR to medicine, chemistry, psychology, and economics, showing that a single cross-domain reward model can supervise all four areas without task-specific tweaks Su et al. (2025a). Focusing on healthcare, Med-RLVR applies the same recipe to multiple-choice clinical QA, gaining eight accuracy points over supervised baselines while eliciting chain-of-thought reasoning from a 3B-parameter base model Zhang et al. (2025). These studies collectively show promising progress in taking RL beyond math and code, yet sparse reward signals, unreliable verifiers, and limited benchmark coverage remain open challenges.
Rubric-based Evaluation
Concise, task-specific rubrics are now standard for evaluating frontier LLMs on open-ended benchmarks (Arora et al., 2025; Ruan et al., 2025; Hashemi et al., 2024; Pathak et al., 2025). Pathak et al. show that rubric-prompted GPT-4 graders are more accurate and consistent than a single question-agnostic checklist (Pathak et al., 2025), yet their rubrics appear only at evaluation time, not in training. HealthBench extends the idea to medicine, pairing 48 k clinician-written criteria with GPT-4 judges to score factuality, safety, and empathy (Arora et al., 2025). Moving from evaluation to tuning, Configurable Preference Tuning (CPT) synthesizes rubric-conditioned preference pairs for DPO fine-tuning (Gallego, 2025). None of these approaches, however, integrates rubrics with verifiable rewards. Rubrics are used only to grade held-out outputs or to generate preference pairs.
Learning from feedback signals
RLHF tunes LLMs with thousands of human preference comparisons, but these subjective labels introduce noise and invite reward hacking (Ouyang et al., 2022). Reinforcement learning with verifiable rewards (RLVR) alleviates these issues by relying on programmatic checks—ranging from exact string matches on GSM8K and MATH to mixed-domain verifiers in General-Reasoner and Cross-Domain RLVR Ma et al. (2025); Su et al. (2025a). Although the resulting signals are low-variance and fully automatic, they can be sparse. Process supervision provides more detailed feedback by rewarding intermediate reasoning steps. MCTS-generated labels and generative reward models such as ThinkPRM improve performance, but at a significant annotation cost (Li et al., 2025; Khalifa et al., 2025). Our rubric based RLVR closes this gap by turning rubric criteria into structured verifiers and using their scalar scores as dense rewards in on-policy RLVR training.
8 Future Work
Interpretability and Reward Robustness
RaR offers a transparent alternative to black-box reward models by explicitly optimizing against human-interpretable rubric criteria. Unlike neural reward models or preference-based methods such as DPO, RaR enables auditability and objective attribution of reward signals. This transparency may also increase robustness to reward hacking, as discrete, interpretable criteria are harder for models to exploit than opaque learned scores.
Future work can explore formal evaluations of RaR’s resilience to reward hacking, including adversarial training scenarios where models attempt to game individual rubric components.
Extending to Real-World Domains and Use Cases
Rubric-guided post-training has been underexplored in open-ended or agentic tasks. A promising direction is expanding RaR to domains with subjective or multi-step goals such as tool use, reasoning under uncertainty, or real-world decision support-where rubrics can provide structured, task-specific guidance.
Rubrics as Curriculum
Rubrics naturally encode hierarchical task structure—from essential facts to optional stylistic preferences—which can support implicit curriculum learning. Early in training, models may only satisfy simpler criteria; as learning progresses, more complex targets become attainable. Future work can investigate dynamic weighting or staged introduction of rubric items to further leverage this curriculum effect.
9 Conclusion
We introduced Rubrics as Rewards (RaR), a framework for post-training language models using structured, checklist-style rubrics as reward signals. By decomposing response evaluation into transparent, multi-criteria objectives—both subjective and objective—RaR provides a modular and interpretable alternative to preference-based methods. Our experiments demonstrate that rubric-guided training achieves strong performance across domains, significantly outperforming Likert-based baselines and matching or exceeding the performance of reference-based reward generation. Additionally, we show that RaR improves alignment with human preferences while reducing dependence on large judge models.
Limitations
Our study is limited to two domains (medical and science), and further validation is needed to assess generalizability to other tasks, such as open-ended dialogue. We explore only two reward aggregation strategies—implicit and explicit—and do not investigate alternative weighting schemes. Additionally, we do not conduct controlled analyses of reward hacking risks. Finally, our judges are off-the-shelf LLMs; future work could explore dedicated evaluators with stronger reasoning capabilities.
Acknowledgments
We thank Qin Lyu, Nikhil Barhate, and Zijian Hu for supporting this research through the development of the in-house post-training platform RLXF. We also thank Robert Vacareanu for valuable early feedback and discussions.
Appendix A Training dataset breakdown
In this section, we show the topic distribution breakdown of the medicine and science training dataset and the rubric type distribution breakdown for the corresponding dataset.
A.1 Medicine training dataset - Overall Data Distribution
Metric Value
Total examples 20 166
Avg. rubrics per question 7.5
Avg. question length (words) 45.0
Table 5: Aggregate statistics for the full medical (train and validation) dataset.
Rubric Type Count Percent
Important 52 748 34.1
Essential 47 584 30.7
Optional 34 261 22.1
Pitfall 20 215 13.1
Table 6: Rubric-type distribution across all 20,166 examples.
Topics Count Percent
Total examples 20 166 100.0
Medical Diagnosis 10 147 50.3
Medical Treatment 3 235 16.0
Medical Knowledge 2 557 12.7
Medical Diag. and Mngmnt 2 033 10.1
Medical Biology 770 3.8
Other 428 2.1
Medical Ethics 377 1.9
Health Physics 276 1.4
Epidemiology & Pub. Health 216 1.1
General Medicine 113 0.6
Forensic Medicine 14 0.1
Table 7: Distribution of topics in the medical training and validation dataset
A.2 STEM Dataset - Overall Data Distribution
Metric Value
Total examples 20 625
Avg. rubrics per question 7.5
Avg. question length (words) 52.6
Table 8: Aggregate statistics for the full medical (train and validation) dataset.
Rubric Type Count Percent
Important 52 315 34.8
Essential 42 739 28.4
Optional 33 622 22.3
Pitfall 21 808 14.5
Table 9: Rubric‐type distribution across all 20 625 examples.
Topics Count Percent
Total examples 20625 100.0
General Chemistry 3163 15.3
Quantum Mechanics 3158 15.3
Physical Chemistry 2761 13.4
Statistical Mechanics 2530 12.3
Organic Chemistry 2059 10.0
General Physics 1439 7.0
Condensed Matter Physics 1387 6.7
Genetics 1378 6.7
Molecular Biology 815 4.0
Astrophysics 409 2.0
Inorganic Chemistry 407 2.0
Analytical Chemistry 398 1.9
Electromagnetism 239 1.2
Optics 143 0.7
High Energy Physics 116 0.6
Electromagnetic Theory 105 0.5
Electromagnetics 72 0.3
Relativistic Mechanics 46 0.2
Table 10: Distribution of topics in the STEM training and validation dataset
Appendix B Synthetic Preference Set Generation
We leverage the publicly-released HealthBench Arora et al. (2025) corpus, which contains 5,000 health-related prompts accompanied by expert-written answers. Of these, 4,203 datapoints already include an ideal completion vetted by licensed clinicians. For every such prompt–ideal pair we automatically generate a perturbed counterpart using o3 with the structured template shown below. The template forces the model to (i) spell out a [reasoning] plan for degrading quality, (ii) emit the degraded [perturbed_completion], and (iii) log exact [chunks_added] and [chunks_removed]. Perturbations are accepted only after manual screening confirms that they are objectively worse, along at least one axis of medical accuracy, completeness, clarity, safety, specificity, structure, or tone, while remaining coherent and free of dangerous advice. This procedure produces a balanced evaluation set of 4,203 preferred and 4,203 perturbed responses (8,406 total), which we use in the rubric-versus-Likert experiments in Section 5.
Appendix C Judge Quality impacts on Post-training
We assess whether rubric-guided evaluation improves judge effectiveness compared to rubric-free Likert scoring when used for GRPO training. Table 11 reports judge accuracy on synthetic medical data, with all policies trained using Qwen2.5-7B and varying judge models.
Judge Model RaR-Implicit Simple-Likert
GPT-4o-mini 0.279 0.253
Qwen-32B-Instruct 0.262 0.254
Qwen-14B-Instruct 0.250 0.249
Qwen-7B-Instruct 0.267 0.220
Table 11: Judge quality comparison: rubric-based evaluation vs pure Likert scoring on synthetic medical rubrics.
Rubric-based evaluation consistently yields stronger policies across all judge sizes. The most pronounced improvement appears with Qwen-7B-Instruct (+0.047), where rubric guidance lifts it from weakest to nearly matching larger models. Additionally, rubric-based scores are more tightly clustered (0.250–0.279) than those from Likert-only judges (0.220–0.254), indicating improved consistency.
These results suggest that rubrics help smaller judges approximate high-quality supervision by breaking evaluation into interpretable, binary criteria. This structured approach mitigates scale-related limitations, enabling more reliable reward modeling even with limited-capacity evaluators.
Appendix D Training Details
The training hyperparameters are described in Table 12.
Hyperparameters
num_rollouts_per_prompt 16
batch_size (effective) 96
sampling_temperature 1.0
warmup_ratio 0.1
learning_rate 5.0e-06
lr_scheduler_type constant_with_warmup
max_length 3584
num_train_steps 300
Table 12: GRPO hyperparameter settings for Medical and Science domains.
Appendix E LLM-Judge Prompts
Prompt for RUBRIC-LIKERT-IMPLICIT Method System Prompt: You are an expert evaluator. Given a user prompt, a generated response, and a list of quality rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics. Consider all rubrics holistically when determining your score. A response that violates multiple rubrics should receive a lower score, while a response that satisfies all rubrics should receive a higher score. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json
"rating": 7
User Prompt Template: Given the following prompt, response, and rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics. <prompt> prompt </prompt> <response> response </response> <rubrics> rubric_list_string </rubrics> Your JSON Evaluation:
Prompt for SIMPLE-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt and a generated response, please rate the overall quality of the response on a scale of 1 to 10, where 1 is very poor and 10 is excellent. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json
"rating": 8
User Prompt Template: Given the following prompt, and response, please rate the overall quality of the response on a scale of 1 to 10. <prompt> prompt </prompt> <response> response </response> Your JSON Evaluation:
Prompt for REFERENCE-LIKERT Baseline System Prompt: You are an expert evaluator. Given a user prompt, a reference response, and a generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference response. Consider factors such as accuracy, completeness, coherence, and helpfulness when comparing to the reference. The reference response represents a high-quality answer that you should use as a benchmark. Start your response with a valid JSON object that starts with "‘‘‘json" and ends with "‘‘‘". The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10. Example response: ‘‘‘json { "rating": 8 }‘‘‘ User Prompt Template: Given the following prompt, reference response, and generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference. <prompt> prompt </prompt> <reference_response> reference </reference_response> <generated_response> response </generated_response> Your JSON Evaluation:
Appendix F Evaluation Prompts
GPQA Evaluation Prompt Determine whether the following model response matches the ground truth answer. ## Ground truth answer## Option correct_answer or correct_answer_text ## Model Response ##: response_text A response is considered correct if it’s final answer is the correct option letter (A, B, C, or D), or has the correct answer text. Please respond with only "Yes" or "No" (without quotes). Do not include a rationale.
Appendix G Predefined Rubrics
Predefined Rubrics for RaR-Predefined Method The response contains correct information without factual errors, inaccuracies, or hallucinations that could mislead the user. The response fully answers all essential parts of the question and provides sufficient detail where needed. The response is concise and to the point, avoiding unnecessary verbosity or repetition. The response effectively meets the user’s practical needs, provides actionable information, and is genuinely helpful for their situation.
Appendix H Synthetic Rubric Generation Prompts
Prompt for Synthetic Rubrics Generation: Medical Domain You are an expert rubric writer. Your job is to generate a self-contained set of evaluation criteria (“rubrics”) for judging how good a response is to a given question. Rubrics can cover aspects of a response such as, but not limited to, factual correctness, ideal-response characteristics, style, completeness, helpfulness, harmlessness, patient-centeredness, depth of reasoning, contextual relevance, and empathy. Each item must be self-contained – non expert readers should not need to infer anything or consult external information. Begin each description with its category: “Essential Criteria: …”, “Important Criteria: …”, “Optional Criteria: …”, or “Pitfall Criteria: Does not mention …”. Inputs: question: The full question text. reference_answer: The ideal answer, including any specific facts, explanations, or advice. Total items: Choose 7–20 rubric items based on the complexity of the question. Each rubric item: title (2–4 words). description: One sentence starting with its category prefix that explicitly states exactly what to look for. For example: Essential Criteria: Identifies non-contrast helical CT scan as the most sensitive modality for ureteric stones. Pitfall Criteria: Does not mention identifying (B) as the correct answer. Important Criteria: Explains that non-contrast helical CT detects stones of varying sizes and compositions. Optional Criteria: States “The final answer is (B)” or similar answer choice formatting. weight: For Essential/Important/Optional, use 1–5 (5 = most important); for Pitfall, use –1 or –2. Category guidance: Essential: Critical facts or safety checks; if missing, the response is invalid (weight 5). Important: Key reasoning, completeness, or clarity; strongly affects quality (weight 3–4). Optional: Helpful style or extra depth; nice to have but not deal-breaking (weight 1–2). Pitfall: Common mistakes or omissions specific to this prompt—identify things a respondent often forgets or misstates. Each Pitfall description must begin with “Pitfall Criteria: Does not mention …” or “Pitfall Criteria: Recommends …” and use weight –1 or –2. To ensure self-contained guidance: When referring to answer choices, explicitly say “Identifies (A)”, “Identifies (B)”, etc., rather than vague phrasing. If the format requires a conclusion like “The final answer is (B)”, include a rubric item such as: Essential Criteria: Includes a clear statement “The final answer is (B)”. If reasoning should precede the answer, include a rubric like: Important Criteria: Presents the explanation before stating the final answer. If brevity is valued, include a rubric like: Optional Criteria: Remains concise and avoids unnecessary detail. If the question context demands mention of specific findings, include that explicitly (e.g., “Essential Criteria: Mentions that CT does not require contrast”). Output: Provide a JSON array of rubric objects. Each object must contain exactly three keys—title, description, and weight. Do not copy large blocks of the question or reference_answer into the text. Each description must begin with its category prefix, and no extra keys are allowed. Now, given the question and reference_answer, generate the rubric as described. The reference answer is an ideal response but not necessarily exhaustive; use it only as guidance.
Prompt for Synthetic Rubric Generation: Science Domain You are an expert rubric writer for science questions in the domains of Biology, Physics, and Chemistry. Your job is to generate a self-contained set of evaluation criteria (“rubrics”) for judging how good a response is to a given question in one of these domains. Rubrics can cover aspects such as factual correctness, depth of reasoning, clarity, completeness, style, helpfulness, and common pitfalls. Each rubric item must be fully self-contained so that non-expert readers need not consult any external information. Inputs: question: The full question text. reference_answer: The ideal answer, including any key facts or explanations. Total items: Choose 7–20 rubric items based on question complexity. Each rubric item must include exactly three keys: 1. title (2–4 words) 2. description: One sentence beginning with its category prefix, explicitly stating what to look for. For example: Essential Criteria: States that in the described closed system, the total mechanical energy (kinetic plus potential) before the event equals the total mechanical energy after the event. Important Criteria: Breaks down numerical energy values for each stage, demonstrating that initial kinetic energy plus initial potential energy equals final kinetic energy plus final potential energy. Optional Criteria: Provides a concrete example, such as a pendulum converting between kinetic and potential energy, to illustrate how energy shifts within the system. Pitfall Criteria: Does not mention that frictional or air-resistance losses are assumed negligible when applying conservation of mechanical energy. 3. weight: For Essential/Important/Optional, use 1–5 (5 = most important); for Pitfall, use –1 or –2. Category guidance: Essential: Critical facts or safety checks; omission invalidates the response. Important: Key reasoning or completeness; strongly affects quality. Optional: Nice-to-have style or extra depth. Pitfall: Common mistakes or omissions; highlight things often missed. Format notes: When referring to answer choices, explicitly say “Identifies (A)”, “Identifies (B)”, etc. If a clear conclusion is required (e.g. “The final answer is (B)”), include an Essential Criteria for it. If reasoning should precede the final answer, include an Important Criteria to that effect. If brevity is valued, include an Optional Criteria about conciseness. Output: Provide a JSON array of rubric objects. Each object must contain exactly three keys—title, description, and weight. Do not copy large blocks of the question or reference_answer into the text. Each description must begin with its category prefix, and no extra keys are allowed. Now, given the question and reference_answer, generate the rubric as described. The reference answer is an ideal response but not necessarily exhaustive; use it only as guidance.
Appendix I Perturbed Dataset Generation
Prompt for Health Response Perturbation You are tasked with creating perturbed versions of health-related responses to generate training data for evaluating model robustness. Given an original health prompt and an ideal completion, you must generate a perturbed version that is strictly worse than the ideal response but still plausible. Input Format: [prompt]: prompt [ideal_completion]: ideal_completion Task Instructions: For the perturbed version, follow this exact process: [reasoning]: Explain your strategy for making the response worse. Consider these perturbation approaches: - Medical Accuracy: Introduce subtle medical inaccuracies or outdated information - Completeness: Remove crucial information or context - Clarity: Make explanations less clear or more confusing - Safety: Reduce appropriate medical disclaimers or caveats - Specificity: Replace specific advice with vague generalizations - Structure: Disorganize the logical flow of information - Tone: Make the response less empathetic or more dismissive [perturbed_completion]: Create the modified response based on your reasoning strategy. [chunks_added]: List any new text segments you added (even if they’re incorrect or unhelpful) [chunks_removed]: List any text segments you removed from the ideal response Quality Guidelines: - Ensure Degradation: Each perturbed response must be objectively worse than the ideal response - Maintain Plausibility: The perturbed responses should still seem like reasonable attempts (not obviously nonsensical) - Vary Approaches: Use different perturbation strategies for each version - Track Changes Precisely: Be specific about what text was added or removed - Health Context Sensitivity: Be mindful that health misinformation can be harmful - focus on subtle degradations rather than dangerous advice Example Perturbation Strategies: - Remove important safety warnings or medical disclaimers - Add outdated medical information that was once accepted but is now known to be incorrect - Remove specific dosage information or timing details - Add overly general statements that replace specific guidance - Remove context about when to seek professional medical help - Add confusing or contradictory information - Remove step-by-step instructions and replace with vague advice - Add unnecessarily complex medical jargon without explanation

### v2

@@
 Abstract
-Extending Reinforcement Learning with Verifiable Rewards (RLVR) to real-world tasks often requires balancing objective and subjective evaluation criteria.
-However, many such tasks lack a single, unambiguous ground truth—making it difficult to define reliable reward signals for post-training language models.
-While traditional preference-based methods offer a workaround, they rely on opaque reward functions that are difficult to interpret and prone to spurious correlations.
-We introduce Rubrics as Rewards (RaR), a framework that uses structured, checklist-style rubrics as interpretable reward signals for on-policy training with GRPO.
-Our best RaR method yields up to a 28% relative improvement on HealthBench-1k compared to simple Likert-based approaches, while matching or surpassing the performance of reward signals derived from expert-written references.
-By treating rubrics as structured reward signals, we show that RaR enables smaller-scale judge models to better align with human preferences and sustain robust performance across model scales.
-Rubrics as Rewards:
-Reinforcement Learning Beyond Verifiable Domains
-Anisha Gunjal   Anthony Wangthanks: Work done while at Scale AI.   Elaine Lau Vaskar NathBing LiuSean Hendryx Scale AI anisha.gunjal@scale.com
+Reinforcement Learning with Verifiable Rewards (RLVR) has proven effective for complex reasoning tasks with clear correctness signals such as math and coding.
+However, extending it to real-world reasoning tasks is challenging, as evaluation depends on nuanced, multi-criteria judgments rather than binary correctness.
+Instance-specific rubrics have recently been used in evaluation benchmarks to capture such judgments, but their potential as reward signals for on-policy post-training remains underexplored.
+We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning method that extends RLVR beyond verifiable domains by using rubric-based feedback.
+Across both medical and science domains, we evaluate multiple strategies for aggregating rubric feedback into rewards.
+The best RaR variant achieves relative improvements of up to 31% on HealthBench and 7% on GPQA-Diamond over popular LLM-as-judge baselines that rely on direct Likert-based rewards.
+These results demonstrate that RaR-trained policies adapt well to diverse evaluation formats, performing strongly on both rubric-based and multiple-choice tasks.
+Moreover, we find that using rubrics as structured reward signals yields better alignment for smaller judges and reduces performance variance across judge scales.
+# anisha.gunjal@scale.com
 1 Introduction
-Reinforcement Learning with Verifiable Rewards (RLVR) has enabled large language models to elicit complex reasoning on tasks with verifiable outcomes, most prominently in mathematics and coding, where a reward model can be replaced with a scoring function and final answers are easily verifiable or test cases provide automated mechanisms for evaluation Lambert et al.
-(2024); Guo et al.
-(2025a); Cui et al.
-(2025).
-However, many real-world tasks lack such explicit verifiable answers, leaving models without direct reward feedback.
-In practice, researchers often turn to RLHF via preference ranking—collecting human judgments over pairs or lists of model outputs to fill this gap.
-While preference-based reward models can bootstrap performance, they tend to overfit superficial artifacts (e.g.
-response length, formatting quirks, annotator biases) Singhal et al.
-(2023); Wang et al.
-(2024); Chen et al.
-(2024b); Ye et al.
-(2024); Gudibande et al.
-(2023) and require large volumes of pairwise comparisons Ouyang et al.
-(2022), making preference-based reward models brittle and costly.
-To address this gap, we introduce Rubrics as Rewards (RaR) for on-policy Reinforcement Learning that treats structured criteria or rubrics as the core reward mechanism.
-We use rubrics as checklists Arora et al.
-(2025); Sirdeshmukh et al.
-(2025) composed of independent subgoals allowing for automatable feedback aligned with expert intent.
+Reinforcement Learning with Verifiable Rewards (RLVR) has enabled large language models to elicit complex reasoning on tasks with clear verifiable outcomes.
+This is especially effective in domains like math and code, where reward models can be replaced by scoring functions or test cases that automatically verify correctness [19, 11, 6].
+However, extending RLVR to unstructured, real-world reasoning is challenging because such tasks lack easily verifiable answers.
+A common workaround is to use preference-based reward models, but they tend to overfit superficial artifacts (e.g.
+response length, formatting quirks, annotator biases) [29, 34, 4, 36, 10] and require large volumes of pairwise comparisons [25].
+Instance-specific rubrics have recently emerged for nuanced evaluation in expert domains [2], yet their application in on-policy training for expert-level reasoning is largely unexplored.
+To address this gap, we explore a paradigm shift that introduces a middle ground between the simplicity of verifiable rewards and the expressiveness of preference rankings, which often come with human artifacts and operational overhead.
+We introduce Rubrics as Rewards (RaR), a framework for on-policy Reinforcement Learning that uses structured criteria or rubrics as the core reward mechanism.
+Rather than using rubrics only for evaluation [2, 30], we treat them as checklist-style supervision that produces reward signals for on-policy RL.
+Each rubric is composed of modular, interpretable subgoals that provide automatable feedback aligned with expert intent.
 By decomposing “what makes a good response” into tangible, human-interpretable criteria, rubrics offer a middle ground between binary correctness signals and coarse preference rankings.
@@
 (ii) GRPO Training: These rubrics are used to prompt an LLM judge for reward estimation, which drives policy optimization via the GRPO on-policy learning loop.
-Previous works train generative reward models that learn to evaluate reasoning or final outputs with interpretable scores Chen et al.
-(2025); Whitehouse et al.
-(2025); Anugraha et al.
-(2025); Guo et al.
-(2025b), and some have even used a model’s internal confidence estimates as a proxy for reward Zhao et al.
-(2025).
-Concurrently, recent efforts have extended verifiable datasets beyond STEM domains, broadening the applicability of RLVR methods to a wider range of tasks Su et al.
-(2025b); Ma et al.
-(2025).
-However, developing a general-purpose approach for specifying reliable and scalable rewards remains an open challenge, particularly in settings where no single ground truth exists, or where both subjective and objective criteria must be considered to evaluate response quality.
-The Rubrics as Rewards strategy offers a flexible solution by repurposing structured evaluation criteria into multi-dimensional reward signals.
-Figure 1 illustrates our approach for generating rubrics and leveraging them as reward signals for on-policy training.
-Our key contributions are as follows: (i) We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning framework that uses checklist-style rubrics to supervise multi-criteria tasks, enabling stable training and improved performance in both reasoning and real-world domains.
-(ii) We apply our rubric generation approach to the medicine and science domains, producing two training datasets—RaR-Medicine-20k and RaR-Science-20k [^1] The datasets will be released soon..
-(iii) We show that models trained with RaR match or outperform strong baselines, achieving notable gains in accuracy across diverse domains.
-(iv) By using rubrics as structured reward signals, RaR enables smaller judge models to achieve better alignment with human preferences and maintains robust performance across different model scales.
+Previous works train generative reward models that learn to evaluate reasoning or final outputs with interpretable scores [5, 35, 1, 12], and some have even used a model’s internal confidence estimates as a proxy for reward [39].
+More recent efforts have extended verifiable datasets beyond STEM domains, broadening the applicability of RLVR methods to a wider range of tasks [32, 23].
+Yet a general-purpose approach for specifying reliable reward signals remains elusive, particularly in tasks without a single ground truth where both subjective and objective criteria must be considered.
+In contrast, we treat rubrics as instance-specific, reusable reward functions.
+Once generated, rubrics provide interpretable and automatable supervision that can be applied consistently across new rollouts, offering a scalable and transparent alternative to opaque reward modeling in on-policy learning.
+Recent concurrent works explore checklists and principled rubric criteria for preference tuning and LLM safety [8, 33, 7], highlighting a growing trend toward structured supervision.
+In contrast, we convert rubrics into reward functions for on-policy RL, targeting expert reasoning and applied real-world domains.
+This closes the rubric-to-learning loop and improves performance on both rubric-guided evaluations and tasks with verifiable answers.
+Figure 1 illustrates our framework.
+Our key contributions are as follows: (i) We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning framework that uses checklist-style rubrics for multi-criteria supervision in reasoning and real-world domains.
+(ii) We synthesize instance-specific rubrics for medicine and science and release the corresponding training sets, RaR-Medicine  [^1] RaR-Medicine: https://huggingface.co/datasets/anisha2102/RaR-Medicine and RaR-Science  [^2] RaR-Science: https://huggingface.co/datasets/anisha2102/RaR-Science.
+(iii) RaR-trained models consistently outperform strong baselines and yield a stable, generalizable training signal, with gains on both rubric-scored and verifiable multiple-choice evaluation settings.
+(iv) Our results demonstrate that rubric-based rewards provide stable supervision across judge sizes, helping smaller models align effectively with human preferences and maintaining robust evaluation performance from small to large judges.
 2 Rubrics as Rewards
@@
 Let $x$ denote an input prompt and $\hat{y}\sim\pi_{\theta}(\cdot\mid x)$ be a sampled response from a model parameterized by $\theta$.
-In domains without single ground-truth answers or automatic correctness signals, we define a structured reward function using prompt-specific rubric criteria.
+In domains without single ground-truth answers or automatic correctness signals, we define a structured reward function using instance-specific rubric criteria.
 Each prompt $x$ is associated with a set of $k$ rubric items $\{(w_{j},c_{j})\}_{j=1}^{k}$, where $w_{j}\in\mathbb{R}$ denotes the weight of criterion $j$, and $c_{j}:(x,\hat{y})\mapsto\{0,1\}$ is a binary correctness function that indicates whether the response $\hat{y}$ satisfies that criterion given the prompt.
-The final normalized scalar reward is computed as:
+2.2 Reward Aggregation Strategies
+We investigate two complementary approaches for combining rubric feedback into scalar rewards:
+Explicit Aggregation.
+Each criterion is independently evaluated using an LLM-as-judge, and the final normalized reward is computed as:
 $r(x,\hat{y})=\frac{\sum_{j=1}^{k}w_{j}\cdot c_{j}(x,\hat{y})}{\sum_{j=1}^{k}w_{j}}$ (1)
-This normalization ensures reward values are consistent across prompts with varying numbers and weights of rubric items.
-Explicit Rubric Aggregation.
-Each criterion is evaluated independently using a generative reward model, and the reward is computed using Equation 1.
-While we currently assume binary values for $c_{j}$, this framework supports extension to continuous-valued judgments.
-Implicit Rubric Aggregation.
-An alternative approach passes all rubric descriptions and weights to a generative reward model, which internally computes a scalar reward:
-$r_{\text{implicit}}(x,\hat{y})=f_{\phi}(x,\hat{y},\{(w_{j},d_{j})\}_{j=1}^{k})$ (2)
-Here, $f_{\phi}$ denotes an LLM-based judge that takes as input the prompt $x$, the generated response $\hat{y}$, and a set of weighted rubric criteria.
-Each pair $(w_{j},d_{j})$ consists of a scalar or categorical weights $w_{j}\in\mathbb{R}$ and a rubric criterion $d_{j}$ describing a specific aspect of response quality.
-This formulation delegates the aggregation of evaluation criteria to the model itself, allowing it to compute a holistic reward score.
-In practice, we use an LLM-as-judge with separate prompt templates containing prompt, response, and rubric items for explicit and implicit aggregation.
-The prompts used in our experiments are detailed in Appendix E.
+Normalization makes rewards comparable across prompts that differ in rubric count or weights.
+Although we use binary checks for $c_{j}$ in our experiments, the formulation can be extended to continuous-valued scores.
+Implicit Aggregation.
+All rubric criteria along with categorical weights are passed to an LLM-as-judge, delegating the aggregation to the model itself to produce a single scalar reward:
+$r_{\text{implicit}}(x,\hat{y})=f_{\phi}(x,\hat{y},\{d_{j}\}_{j=1}^{k})$ (2)
+Here, $f_{\phi}$ denotes an LLM-based judge that takes the prompt $x$, the response $\hat{y}$, and the set of rubric criteria $\{d_{j}\}$ as input.
+This formulation allows the model to compute a holistic reward score directly, avoiding the need to manually tune rubric weights.
+The prompts used for each method are detailed in Appendix A.6.
 2.3 Generalization of RLVR with Rubrics as Rewards
@@
 3.1 Desiderata
-We synthesize prompt-specific rubrics guided by a set of design principles.
-Each rubric defines the criteria for a high-quality response and enables human-interpretable supervision.
-We use large language models (LLMs) as expert proxies to generate these rubrics while ensuring adherence to the following desiderata:
-Grounded in Expert-guidance
-Reference answers produced by human experts or stronger LLMs serve as proxies for expert guidance.
-Rubrics are grounded in these references to capture the key facts, reasoning steps, and conclusions necessary for correctness.
-Comprehensive coverage
-Rubrics are designed to span multiple quality dimensions, including factual accuracy, logical structure, completeness, style, and common pitfalls.
-Negative “pitfall” criteria help penalize frequent or high-risk errors.
-The total number of rubric items per rubric (typically 7–20) is chosen to balance coverage with signal sparsity.
-Semantic weighting
-Each criterion is labeled with a categorical importance level (e.g., Essential, Important, Optional, Pitfall) that reflects its relative priority in the final reward.
-This supports interpretable reward aggregation.
-Future work may explore numeric or learned weighting for greater flexibility.
-Self-contained evaluation
-Rubrics are written to be independently actionable-each item can be evaluated in isolation by either human annotators or LLM-based judges, without requiring external context or domain knowledge.
-3.2 Synthesis
-We apply the above framework on real world science and Medicine reasoning domains.
-For each domain, the prompt (included in Appendix H) instructs the LLM to generate 7–20 rubric items based on the complexity of the input question.
-Each item is assigned a categorical weight (e.g., Essential Criteria, Important Criteria) to determine its importance to a correct answer.
-The rubrics are designed to be fully self-contained which means that non-expert readers should be able to evaluate response quality using only the rubric.
-The evaluation outcome can be directly used as reward functions during reinforcement learning.
-We generate synthetic rubrics using OpenAI’s o3-mini and gpt-4o models OpenAI o3-mini (2025); Jaech et al.
-(2024); Hurst et al.
-(2024) as described in Section 3.
-The RaR-Medical-20k and RaR-Science-20k are released for public use.
-For both domains, we use the reference answers provided in the datasets as a proxy for human expertise while generating synthetic rubrics.
-These rubric sets are then used to supervise the training of smaller models through either explicit or implicit reward aggregation (Section 2.1).
+A rubric specifies criteria for high-quality responses and provides human-interpretable supervision. We identify four desiderata for effective rubric generation:
+Grounded in Expert Guidance.
+Rubrics should reflect domain expertise by capturing the essential facts, reasoning steps, and conclusions necessary for correctness.
+Ideally, this grounding comes from human experts or their high-quality proxies.
+Comprehensive Coverage.
+Rubrics should span multiple dimensions of response quality, including factual accuracy, logical coherence, completeness, style, and safety.
+Negative criteria (pitfalls) help identify frequent or high-risk errors that undermine overall quality.
+Criterion Importance.
+Rubrics should reflect that some dimensions of response quality are more critical than others.
+For example, factual correctness must outweigh secondary aspects such as stylistic clarity.
+Assigning weights to criteria ensures this prioritization, whether through simple categorical tags, explicit numeric values, or learned weighting schemes.
+Self-Contained Evaluation.
+Each rubric item should be independently actionable, allowing either human annotators or automated judges to assess it in isolation without requiring external context or domain-specific knowledge.
+3.2 Rubrics Creation
+We apply these desiderata to datasets for reasoning tasks in medicine and science.
+Given the scarcity of human-annotated rubric datasets in these domains, we use LLMs to generate instance-specific rubrics from golden reference answers at scale, enabling the study of structured rewards without costly human annotation.
+For each prompt, an LLM generates a rubric of 7–20 self-contained items.
+Each item is assigned both a numeric and a categorical weight reflecting its relative importance.
+While numeric weights provide fine-grained prioritization, in our experiments we adopt categorical labels (Essential, Important, Optional, Pitfall) for ease of implementation and interpretability in controlled settings.
+The resulting rubrics are then used directly as reward functions through either explicit aggregation (Eq.
+1) or implicit aggregation (Sec.
+2.2).
+In practice, we generate rubrics using OpenAI’s o3-mini and GPT-4o [24, 15, 14], conditioning generation on reference answers from the underlying datasets to approximate expert grounding.
+The resulting collections—RaR-Medicine and RaR-Science—are released for public use.
+These rubric sets supervise smaller policies under GRPO using both explicit and implicit reward aggregation.
 4 Experiments
-Domain $\rightarrow$ Medicine Science
-Method/Baseline $\downarrow$ HealthBench-1k GPQA_Diamond
-Qwen2.5-7b 0.0818 0.3030 $\pm$ 0.0286
-Qwen2.5-7b-Instruct 0.2359 0.3598 $\pm$ 0.0077
-Simple-Likert 0.2489 0.3409 $\pm$ 0.0104
-Reference-Likert 0.3155 0.3775 $\pm$ 0.0350
-Predefined RaR 0.2472 0.3485 $\pm$ 0.0365
-RaR-Explicit (o3-mini rubrics) 0.2559 0.3333 $\pm$ 0.0504
-RaR-Explicit (GPT-4o rubrics) 0.2979 0.3030 $\pm$ 0.0197
-RaR-Implicit (o3-mini rubrics) 0.3107 0.3864 $\pm$ 0.0407
-RaR-Implicit (GPT-4o rubrics) 0.3194 0.3662 $\pm$ 0.0191
-Table 1: Performance of baseline and RaR (Rubrics as Rewards) methods on medicine and science domains.
-For HealthBench-1k, we report the overall score; for GPQA_Diamond, we report mean accuracy over 4 runs with standard deviation.
-All experiments use gpt-4o-mini as the judge model.
-(RaR) methods are trained using rubrics generated by either o3-mini or GPT-4o.
-Across both domains, the RaR-Implicit method consistently outperforms Simple-Likert and matches or exceeds the performance of Reference-Likert.
 4.1 Datasets
-We investigate the utility of rubrics as rewards across two reasoning domains medicine and science.
-  • RaR‑Medical‑20k: We curate a set of 20k prompts from medicine‑related subsets of medical‑o1‑reasoning‑SFT Chen et al.
-(2024a), the natural_reasoning dataset Yuan et al.
-(2025), the SCP‑116K dataset Lu et al.
-(2025), and the GeneralThought‑430K dataset General Reasoning (2025).
-  • RaR‑Science‑20k: We curate roughly 20k science prompts aligned with the GPQA Diamond benchmark from the natural_reasoning dataset Yuan et al.
-(2025), the SCP‑116K dataset Lu et al.
-(2025), and the GeneralThought‑430K dataset General Reasoning (2025).
-Refer Tables 7 and 10 in Appendix for the full topic distributions.
+We investigate the utility of rubrics as rewards across two reasoning domains, medicine and science.
+  • RaR-Medicine: A dataset of 20k prompts drawn from diverse medical reasoning sources, including medical-o1-reasoning-SFT [3], natural_reasoning [37], SCP-116K [22], and GeneralThought-430K [9].
+Instance-specific rubrics for this dataset are generated with GPT-4o (see Appendix A.1).
+  • RaR-Science: A dataset of $\sim$20k prompts curated to align with GPQA-Diamond categories.
+Prompts are sourced from natural_reasoning [37], SCP-116K [22], and GeneralThought-430K [9], covering a broad range of scientific reasoning tasks (Appendix A.2).
+Rubrics for this dataset are synthesized with o3-mini.
 4.2 Training Details
-We train all models using the GRPO algorithm Shao et al.
-(2024), with Qwen2.5-7B as the base policy.
-Training is conducted with a batch size of 96, a learning rate of $5\times 10^{-6}$, and a constant learning rate schedule with 10% linear warmup.
-Full hyperparameter details are provided in Appendix 11.
-All models are trained on a single compute node equipped with 8 NVIDIA H100 GPUs.
+We conduct all experiments using on-policy reinforcement learning with the GRPO algorithm [28], taking Qwen2.5-7B as the base policy.
+Models are trained with a batch size of 96, a learning rate of $5\times 10^{-6}$, and a constant schedule with 10% linear warmup.
+Complete hyperparameter settings are listed in Appendix 11.
+Training runs are executed on a single compute node equipped with 8 NVIDIA H100 GPUs.
 Our training pipeline consists of the following key components:
 Response Generation: For each prompt $q$, we sample $k=16$ responses from the current policy $\pi_{\theta}$, using a context length of 3584 and a sampling temperature of 1.0.
-Reward Computation: We use gpt-4o-mini as the judge model to assign rewards $R_{q}$ to the sampled responses. The reward functions are described in Sections 4.3 and 4.4.
-Policy Update: The policy is updated using GRPO based on the computed rewards.
-4.3 Baselines
-We consider compare various baselines and off-the-shelf post-trained models. All rubric-free baselines are with Qwen2.5-7b as the base policy.
-For off-the-shelf baselines we evaluate performance on Qwen2.5-7b. We also include the performance of Qwen2.5-7b-Instruct to compare with instruction tuned variant of the base policy.
-We describe the post-trained baselines without rubrics below:
-SIMPLE-LIKERT:
-LLM-judge is asked to output a Likert Score from 1-10 (normalized to 0-1) for each response-prompt pair.
+Reward Computation with Rubrics: We use gpt-4o-mini as the judge model to assign rewards $R_{q}$ to the sampled responses.
+We experiment with various reward computation and aggregations strategies further described in Sections 4.3 and 4.4.
+Policy Update: The policy weights are updated using GRPO based on the computed rewards.
+4.3 Rubric-Free Baselines
+We consider various rubric-free baselines and off-the-shelf post-trained models. Rubric-free baselines are trained with Qwen2.5-7B as the base policy.
+OFF-THE-SHELF:
+For off-the-shelf baselines we evaluate performance on Qwen2.5-7B. We also include the performance of Qwen2.5-7B-Instruct to compare with instruction-tuned variant of the base policy.
+DIRECT-LIKERT:
+An LLM-as-judge provides a direct assessment for each response–prompt pair on a 1–10 Likert scale [40, 17], normalized to $[0,1]$.
+The resulting score is used directly as the reward signal for training.
 REFERENCE-LIKERT:
-The judge model compares the generated response against a high-quality reference answer and outputs a Likert score from 1–10, which is then normalized to the $[0,1]$ range.
+An LLM-as-judge compares the generated response against a reference answer (written by experts or stronger LLMs) and assigns a 1–10 Likert score [40], normalized to $[0,1]$.
+This reference-guided score is used as the reward signal for policy updates.
 The reward for each (prompt, response, reference) triplet is defined as:
 $R_{\text{ref}}(q,x)=\mathrm{Norm}(\text{LikertScore}(q,x,x^{*}))$
-where $x^{*}$ denotes the reference answer, either written by a human or generated by a stronger LLM.
+where $x^{*}$ denotes the reference answer.
 4.4 Rubric-guided Methods
-PREDEFINED-RaR:
-This method uses a fixed set of generic rubrics (not prompt-specific, as described in Appendix C) and computes the reward via an explicit weighted sum, following Equation 1.
-All rubric criteria are weighted uniformly.
+RaR-PREDEFINED:
+This method uses a fixed set of generic rubrics for all prompts (e.g.
+response is concise, response contains correct information).
+It employs the Explicit Aggregation method (Equation 1) with all criteria weighted uniformly (see Appendix A.5).
 RaR-EXPLICIT:
-This variant also employs explicit aggregation (1) but uses prompt-specific rubrics generated by the method in Section 3.
-Since the generated rubrics include categorical labels, we assign numerical weights using a simple default scheme: {"Essential": 1.0, "Important": 0.7, "Optional": 0.3, "Pitfall": 0.8}, with higher penalty applied to violated pitfalls.
-[^2] Pitfall criteria are phrased in positive form (e.g., “The response avoids misinformation”), so satisfying them contributes positively to the score.
-If a pitfall is not satisfied, the corresponding reward is reduced or penalized.
+This variant also uses Explicit Aggregation using a weighted sum (Equation 1) but applies it to instance-specific rubrics from Section 3.
+We manually assign numerical weights based on the generated categorical labels: {"Essential": 1.0, "Important": 0.7, "Optional": 0.3, "Pitfall": 0.9} [^3] Pitfall criteria are phrased in positive form (e.g., “The response avoids misinformation”), so satisfying them contributes positively to the score.
+If a pitfall is not satisfied, the corresponding reward is reduced or penalized..
 RaR-IMPLICIT:
-This variant also leverages prompt-specific rubrics; however, instead of aggregating criterion scores explicitly, it prompts the judge model to evaluate the response holistically and assign a single Likert rating (1–10).
-This rating is then normalized to the $[0,1]$ range:
-$R_{\text{implicit}}(q,x)=\mathrm{Norm}(\text{LikertScore}(q,x,R_{q}))$
-where $R_{q}$ denotes the list of rubrics associated with prompt $q$.
+This variant uses the Implicit Aggregation method (Equation 2).
+It leverages prompt-specific rubrics, where a judge model evaluates the response as a whole to assign a single Likert rating (1–10), avoiding the need for hand-tuned weights.
+The reward is normalized to the $[0,1]$ range during training.
 4.5 Evaluation Setup
-Medical Reasoning
-We evaluate models trained on curated medical reasoning data using HealthBench-1k Arora et al.
-(2025), a 1,000-example subset of HealthBench.
-[^3] We introduce HealthBench-1k as a cost-effective evaluation split for medical reasoning tasks and will release the test set identifiers to support reproducibility..
-We generate model responses with greedy decoding (temperature=0) and report the overall_score.
-Science Reasoning
-We evaluate models trained with curated science set on the GPQA-Diamond subset Rein et al.
-(2024).
-Each evaluation is independently repeated 4 times to account for variance, where each run samples one response per prompt using greedy decoding (temperature=0).
-Answer choices are permuted per example, and model outputs are parsed for boxed answer formats (e.g.,
-boxed{A}).
-If extraction fails, we fall back to a GPT-4o verifier that checks whether the response contains the correct option letter or text (see F).
-Final accuracy is reported as the mean of the 4 independent runs.
+Rubric-Based Evaluation
+We evaluate models trained with RaR-Medicine on HealthBench [2], a benchmark of 5,000 clinical conversations designed to assess model safety and helpfulness in realistic medical scenarios.
+Performance is measured using detailed, physician-authored rubrics.
+We generate responses with greedy decoding (temperature = 0) and report both overall scores and per-axis scores following the original setup.
+For ablation studies, we sample a subset of 1,000 prompts (hereafter referred as HealthBench-1k) and use the rest for training.
+Multiple-Choice Evaluation
+Each model is evaluated across 10 independent runs, using greedy decoding (temperature=0) to sample one response per prompt.
+Answer choices are permuted per example to reduce positional bias, and outputs are parsed for boxed answer formats (e.g., boxed{A}).
+If extraction fails, we fall back to a GPT-4o verifier that checks whether the response contains the correct option letter or text (see Appendix A.4).
+Final accuracy is reported as the mean over 10 runs, and we include 95% confidence intervals to account for run-to-run variance.
+Figure 2: Performance of baselines and RaR (Rubrics as Rewards) variants for the medicine and science domains.
+HealthBench (left): shows per-axis scores across five core axes, with a thin dashed gray line indicating the overall score (all values shown as percentages).
+GPQA-Diamond (right): mean accuracy over 10 runs; error bars represent 95% confidence intervals.
+All policies are evaluated using gpt-4o-mini as the LLM-as-Judge.
+Across both domains, RaR-Implicit consistently outperforms Direct-Likert and demonstrates a competitive advantage over Reference-Likert.
+LLM-Judge Alignment Evaluation
+To measure how well LLM judges align with human preferences, we build a paired evaluation set from roughly 3,000 HealthBench prompts.
+For each prompt, we take the practitioner-approved answer as the preferred response and create a perturbed alternative via controlled edits (see Appendix A.9 for method used for perturbation and prompt selection).
+The metric is pairwise preference accuracy i.e.
+the fraction of pairs where the preferred response scores higher reported across judge models of varying sizes.
 5 Results
 We now present the main findings of our study.
-Rubrics as Rewards Combine Interpretability with Strong Performance.
-Table 1 reports results on the Medicine (HealthBench-1k) and Science (GPQA_Diamond) benchmarks.
-Our method, RaR-Implicit, consistently outperforms baselines such as Simple-Likert, with the best variant achieving upto 28% relative improvement on HealthBench-1k and 13% on GPQA.
-It also surpasses both base and instruction-tuned policy models, demonstrating the effectiveness of rubric-guided training for nuanced response evaluation.
-Furthermore, RaR-Implicit matches or exceeds the performance of the Reference-Likert baseline.
-In our setup, synthetic rubrics are generated using reference answers as proxies for expert supervision, meaning their quality is inherently constrained by the completeness and clarity of those references.
-Despite this, transforming open-ended answers into structured rubric criteria produces reward signals that are both effective and aligned with evaluation objectives.
-We also find that implicit aggregation outperforms explicit, weighted summation which indicates that allowing an LLM-based judge to infer how to balance rubric criteria per prompt is more effective than relying on fixed, hand-crafted weights.
-While explicit aggregation offers more transparency and control, implicit methods better adapt to real-world complexity.
-Rubric-based supervision further enables transparent reward specification by breaking response quality into interpretable criteria.
-Unlike holistic reference comparisons, rubrics reduce stylistic bias and surface-level overfitting.
-Though reference-based Likert scoring may perform slightly better in some cases, rubrics offer a more scalable and controllable approach—particularly where references are scarce or model interpretability is essential.
-Figure 2: LLM Judge Alignment Across Model Scales.
-Rubric-guided evaluation (orange) consistently improves judge accuracy across model sizes compared to pure Likert-based scoring (blue).
-Rubrics without reference answers (green) perform better than Likert baseline, though still lag behind full rubric guidance.
-Smaller scale models benefit the most from rubric structure, narrowing the performance gap with larger judges.
+Rubrics as Rewards shows strong gains across evaluation settings.
+Table 2 reports results on HealthBench (rubric-based, free-form) and GPQA-Diamond (multiple-choice).
+RaR-Implicit consistently outperforms Direct-Likert, with relative gains up to 31% on HealthBench and 7% on GPQA.
+Both rubric-guided variants achieve higher scores than the base and instruction-tuned policies.
+Gains on GPQA-Diamond show that rubric-induced skills generalize beyond rubric-based evaluation.
+The RaR-Predefined variant, which applies a fixed list of generic rubrics to every prompt (no instance-specific synthesis), underperforms because generic criteria miss prompt-specific requirements and common failure modes, producing misaligned reward signals.
+Hence, effective training requires instance-specific rubric synthesis as they better capture task context and typical failure modes.
+Beyond these gains, RaR-Implicit also shows small but consistent gains over Reference-Likert.
+In our setup, rubrics are generated with stronger LLMs using reference answers as proxies for expert supervision, so rubric quality is impacted by reference quality.
+Even so, converting open-ended answers into explicit criteria yields effective, well-aligned reward signals.
+Between the two rubric-guided methods, RaR-Implicit attains the strongest results overall; fixed weighted sums in RaR-Explicit offer more control but can be brittle.
+Explicit weighting can be difficult to tune but offers greater interpretability; we view the choice as application-dependent and leave it to practitioners.
+Future work could explore learned or dynamic weighting strategies that maintain interpretability while improving adaptability.
+Figure 3: Alignment Study of LLM Judges across Model Scales.
+Rubrics as Rewards (orange) consistently improves alignment with human preferences across LLM judge sizes compared to direct Likert-based scoring (blue).
+Judge Alignment using synthetic rubrics without expert grounding (green) outperform the direct Likert baseline, but still fall short of expert-grounded rubrics (orange).
+The rubric structure especially benefits smaller judge models, helping them close the gap with larger models when guided by checklist-style criteria.
 Rubrics enhance alignment with human preferences across model scales
-To evaluate the effectiveness of rubric-based evaluations in reflecting human preferences, we generated datasets consisting preferred (chosen) and perturbed (rejected) responses (see Appendix I for details on dataset generation).
-We use LLM-Judges to assign ratings from 1 to 10 for these responses both with (RaR-IMPLICIT ) and without the guidance of rubrics (SIMPLE-LIKERT).
-Figure 2 shows how often judges correctly rated the preferred responses higher than the rejected ones across various model scales (from 3b to 72b to o3-mini).
-The results consistently show that rubric-guided evaluations provide clearer and more accurate signals, as evidenced by the higher accuracy with which preferred responses receive higher ratings when rubrics are involved.
-This indicates that rubrics improves the alignment with human preferences by offering detailed, context-specific criteria that help judges better discern subtle differences in quality compared to evaluations using only simple Likert-based ratings.
-Further study on LLM-Judge scaling impacts on policy training are discussed in Appendix C.
+We evaluate alignment with humans by having LLM judges of varying sizes score chosen vs.
+rejected HealthBench-1k responses on a 1–10 scale under two settings: (i) rubric-guided (RaR-IMPLICIT), where the instance-specific rubric is provided, and (ii) rubric-free (DIRECT-LIKERT), where only the prompt and answers are shown.
+Figure 3 reports pairwise preference accuracy (the fraction of pairs where the preferred response receives the higher score).
+Rubric guidance improves accuracy for every judge size, with the largest gains for smaller judges, narrowing the gap to larger models.
+This indicates that explicit, context-specific criteria help judges distinguish subtle quality differences better than direct Likert scoring.
+Further analysis of judge-scale effects on GRPO training is detailed in Appendix A.8.
 Expert guidance is crucial for synthetic rubric generation
@@
 The data shows that rubrics developed with reference answers achieve higher accuracy, emphasizing that human insights integrated during rubric generation enable granular criteria and improved alignment with human preferences.
-6 Ablations
-Elements of Rubric Design
-This ablation study examines how the structure and weighting of synthetic rubrics—generated using o3-mini—affect downstream performance on HealthBench.
-As shown in Table 2, predefined generic rubrics substantially underperform compared to prompt-specific ones, underscoring the importance of contextualization.
-Rubrics that include a broader range of criteria—both positive and negative—consistently outperform those limited to essential checks, suggesting that richer evaluation signals lead to better learning.
-Interestingly, we observe minimal performance differences when including rubric weights or pitfall criteria during training.
-One possible explanation is that synthetically generating effective pitfall criteria is inherently difficult, as it requires anticipating the most common or critical failure modes of the model, a task that often demands human intuition and domain expertise.
-As a result, these synthetic negative criteria may lack the specificity or relevance needed to meaningfully penalize undesirable responses.
-Ablation Setting Overall Score
-Predefined Generic Rubrics 0.2898 $\pm$ 0.00526
-Essential-Only Rubrics 0.3562 $\pm$ 0.00976
-No Weights 0.3756 $\pm$ 0.00479
-No-Pitfall Criteria 0.3837 $\pm$ 0.00420
-All Rubrics 0.3723 $\pm$ 0.00005
-Table 2: Ablation results on HealthBench-1k (trained on HealthBench-3.5k subset with Qwen2.5-7B base policy).
-Rubrics generated using o3-mini with access to reference answers.
-Variations test the impact of rubric structure and weighting on implicit reward aggregation.
 Training Method Overall Score
-Expert-Answer-SFT 0.204
-Simple-Likert 0.239
-Reference-Likert 0.317
-RaR-Implicit-Synthetic-NoRef 0.320
-RaR-Implicit-Synthetic 0.359
-RaR-Implicit-Human 0.348
+Expert-Answer-SFT 20.4%
+Simple-Likert 23.9%
+Reference-Likert 31.7%
+RaR-Implicit-Synthetic-NoRef 32.0%
+RaR-Implicit-Synthetic 35.9%
+RaR-Implicit-Human 34.8%
 Table 1: Evaluation on HealthBench: Comparison of human- vs.
@@
 Notably, human-authored rubrics and synthetic rubrics with access to references yield comparable performance.
+Ablation Setting Overall Score
+Essential-Only Rubrics 34.9%
+No Categorical Labels 38.8%
+No Pitfall Criteria 37.2%
+All Rubrics 37.2%
+Table 2: Ablation results for elements of rubric design on HealthBench-1k (trained on HealthBench-3.5k subset with Qwen2.5-7B base policy).
+Rubrics generated using o3-mini with access to reference answers.
+6 Ablations
+Rubric Generation Model Overall Score
+O3-mini (Rubrics with reference) 35.9%
+GPT-4o 34.2%
+GPT-4o-mini 32.7%
+O3-mini 32.4%
+Qwen-72B-Instruct 32.7%
+Qwen-32B-Instruct 31.1%
+Qwen-7B-Instruct 31.9%
+Table 3: Policy performance on HealthBench-1k when trained with GRPO using rubrics generated by different LLMs without reference answers.
+GPT-4o-generated rubrics yield the strongest performance, though they still fall short of rubrics generated with expert (reference-guided) supervision.
+Smaller aligned models (e.g., GPT-4o-mini, O3-mini) remain competitive with larger open-weight models, underscoring the importance of alignment and reasoning ability in rubric generation.
 Impact of Rubric Generation Strategies in Real-World Domains
 How does the method of rubric generation affect downstream training in challenging, real-world settings?
-To investigate this, we evaluate on the HealthBench dataset using a curated subset of 3.5k human-annotated examples, split into train and test sets.
-Table 3 summarizes the results.
-We observe that all rubric-based methods outperform rubric-free baselines, including Reference-Likert.
+To investigate this, we hold out HealthBench-1k for evaluation and use 3.5k prompts from the remaining HealthBench pool to generate rubrics for training as it has access to human generated rubrics.
+Results are summarized in Table 1.
+In-domain testing on HealthBench-1k amplifies RaR’s gains: every instance-specific rubric-based method outperforms rubric-free baselines.
 Notably, even the weakest RaR variant significantly surpasses Reference-Likert, underscoring the advantage of structured supervision in subjective, open-ended domains like healthcare.
@@
 Purely synthetic rubrics, while scalable, currently fall short in capturing the subtle criteria required for robust training in high-stakes domains.
+Elements of Rubric Design
+This ablation study examines how the structure and weighting of synthetic rubrics affect downstream performance on HealthBench-1k.
+As shown in Table 2, rubrics that include a broader range of criteria outperform those limited to essential checks, suggesting that richer evaluation signals lead to better learning.
+Interestingly, we observe minimal performance differences when including rubric weights or pitfall criteria during training.
+One possible explanation is that synthetically generating effective pitfall criteria is inherently difficult, as it requires anticipating the most common or critical failure modes of the model, a task that often demands human intuition and domain expertise.
+As a result, these synthetic negative criteria may lack the specificity or relevance needed to meaningfully penalize undesirable responses.
 Impact of LLM Expertise on Rubric Quality
@@
 Additionally, model attributes such as instruction tuning and reasoning capabilities play a key role in the effectiveness of rubric generation.
-Rubric Generation Model Overall Score
-O3-mini (Rubrics with reference) 0.359
-GPT-4o 0.342
-GPT-4o-mini 0.327
-O3-mini 0.324
-Qwen-72B-Instruct 0.327
-Qwen-32B-Instruct 0.311
-Qwen-7B-Instruct 0.319
-Table 4: Policy performance on HealthBench-1k when trained with GRPO using rubrics generated by different LLMs without reference answers.
-GPT-4o-generated rubrics yield the strongest performance, though they still fall short of rubrics generated with expert (reference-guided) supervision.
-Smaller aligned models (e.g., GPT-4o-mini, O3-mini) remain competitive with larger open-weight models, underscoring the importance of alignment and reasoning ability in rubric generation.
 7 Related Work
 RLVR across domains
-Reinforcement learning with verifiable rewards (RLVR) is quickly extending its reach well beyond math and code.
-General-Reasoner trains on a 200 k mixed corpus-physics, finance, policy, and more-and reports a ten-point jump on MMLU-Pro after GRPO fine-tuning Ma et al.
-(2025).
-A follow-up study extends RLVR to medicine, chemistry, psychology, and economics, showing that a single cross-domain reward model can supervise all four areas without task-specific tweaks Su et al.
-(2025a).
-Focusing on healthcare, Med-RLVR applies the same recipe to multiple-choice clinical QA, gaining eight accuracy points over supervised baselines while eliciting chain-of-thought reasoning from a 3B-parameter base model Zhang et al.
-(2025).
-These studies collectively show promising progress in taking RL beyond math and code, yet sparse reward signals, unreliable verifiers, and limited benchmark coverage remain open challenges.
-Rubric-based Evaluation
-Concise, task-specific rubrics are now standard for evaluating frontier LLMs on open-ended benchmarks (Arora et al., 2025; Ruan et al., 2025; Hashemi et al., 2024; Pathak et al., 2025).
+Reinforcement learning with verifiable rewards (RLVR) is expanding beyond math and code.
+General-Reasoner trains on a 200k mixed corpus spanning physics, finance, and policy, and reports a ten-point gain on MMLU-Pro after GRPO fine-tuning [23].
+A follow-up extends RLVR to medicine, chemistry, psychology, and economics, showing that a single cross-domain reward model can supervise all four without task-specific tweaks [31].
+In healthcare, Med-RLVR applies similar methods to multiple-choice clinical QA, improving accuracy over supervised baselines while eliciting chain-of-thought from a 3B base model [38].
+These results indicate steady progress, yet sparse signals, verifier reliability, and limited benchmark coverage remain open challenges.
+Rubrics for evaluation and training
+Task-specific rubrics are increasingly used to evaluate LLMs in difficult-to-verify domains [2, 27, 13, 26].
 Pathak et al.
-show that rubric-prompted GPT-4 graders are more accurate and consistent than a single question-agnostic checklist (Pathak et al., 2025), yet their rubrics appear only at evaluation time, not in training.
-HealthBench extends the idea to medicine, pairing 48 k clinician-written criteria with GPT-4 judges to score factuality, safety, and empathy (Arora et al., 2025).
-Moving from evaluation to tuning, Configurable Preference Tuning (CPT) synthesizes rubric-conditioned preference pairs for DPO fine-tuning (Gallego, 2025).
-None of these approaches, however, integrates rubrics with verifiable rewards.
-Rubrics are used only to grade held-out outputs or to generate preference pairs.
+show that rubric-prompted LLM graders are more accurate and consistent than a question-agnostic checklist [26].
+HealthBench scales this idea in medicine, pairing 48k clinician-written criteria with GPT-4 judges to score various axes [2].
+Beyond evaluation, rubrics are used to condition preference pairs for DPO (CPT; [8]) and to guide checklist-based preference tuning in safety, instruction-following, and creative-writing settings [33, 7, 18].
+These lines of work primarily use rubrics to grade outputs or to condition preference data, often in non-reasoning domains such as safety, instruction following, or creative writing.
+In contrast, we use rubric criteria directly as reward signals for on-policy RL in expert-reasoning and real-world domains.
 Learning from feedback signals
-RLHF tunes LLMs with thousands of human preference comparisons, but these subjective labels introduce noise and invite reward hacking (Ouyang et al., 2022).
-Reinforcement learning with verifiable rewards (RLVR) alleviates these issues by relying on programmatic checks—ranging from exact string matches on GSM8K and MATH to mixed-domain verifiers in General-Reasoner and Cross-Domain RLVR Ma et al.
-(2025); Su et al.
-(2025a).
-Although the resulting signals are low-variance and fully automatic, they can be sparse.
-Process supervision provides more detailed feedback by rewarding intermediate reasoning steps.
-MCTS-generated labels and generative reward models such as ThinkPRM improve performance, but at a significant annotation cost (Li et al., 2025; Khalifa et al., 2025).
-Our rubric based RLVR closes this gap by turning rubric criteria into structured verifiers and using their scalar scores as dense rewards in on-policy RLVR training.
-8 Future Work
-Interpretability and Reward Robustness
-RaR offers a transparent alternative to black-box reward models by explicitly optimizing against human-interpretable rubric criteria.
-Unlike neural reward models or preference-based methods such as DPO, RaR enables auditability and objective attribution of reward signals.
-This transparency may also increase robustness to reward hacking, as discrete, interpretable criteria are harder for models to exploit than opaque learned scores.
-Future work can explore formal evaluations of RaR’s resilience to reward hacking, including adversarial training scenarios where models attempt to game individual rubric components.
-Extending to Real-World Domains and Use Cases
-Rubric-guided post-training has been underexplored in open-ended or agentic tasks.
-A promising direction is expanding RaR to domains with subjective or multi-step goals such as tool use, reasoning under uncertainty, or real-world decision support-where rubrics can provide structured, task-specific guidance.
-Rubrics as Curriculum
-Rubrics naturally encode hierarchical task structure—from essential facts to optional stylistic preferences—which can support implicit curriculum learning.
-Early in training, models may only satisfy simpler criteria; as learning progresses, more complex targets become attainable.
-Future work can investigate dynamic weighting or staged introduction of rubric items to further leverage this curriculum effect.
+RLHF trains policies with large numbers of human comparisons, which introduces subjectivity and can lead to reward hacking [25].
+RLVR reduces these issues by using programmatic checks, from exact matches on GSM8K and MATH to mixed-domain verifiers in General-Reasoner and Cross-Domain RLVR [23, 31], although signals can be sparse.
+Process supervision [21] provides denser guidance via step-level labels, and MCTS-generated annotations or generative reward models such as ThinkPRM improve performance, but with high annotation cost [20, 16].
+Rubric-based RL finds a middle ground by turning multiple rubric criteria into structured verifiers and using their scalar scores as denser rewards.
 8 Conclusion
@@
 Additionally, we show that RaR improves alignment with human preferences while reducing dependence on large judge models.
-Limitations
-Our study is limited to two domains (medical and science), and further validation is needed to assess generalizability to other tasks, such as open-ended dialogue.
-We explore only two reward aggregation strategies—implicit and explicit—and do not investigate alternative weighting schemes.
-Additionally, we do not conduct controlled analyses of reward hacking risks.
-Finally, our judges are off-the-shelf LLMs; future work could explore dedicated evaluators with stronger reasoning capabilities.
+9 Limitations and Future Work
+Our work focuses on medicine and science to enable controlled experiments.
+This choice allows us to run controlled experiments, but broader validation across dialogue, tool use, or other agentic tasks remains an important direction.
+We evaluate only two reward aggregation strategies, implicit and explicit, since they capture complementary extremes of flexibility and control; future work could explore more advanced ways of combining rubric criteria, such as learning continuous weights for each criterion or dynamically adjusting weights over the course of training to mimic a curriculum (e.g., prioritizing essential correctness early, then gradually emphasizing more subtle qualities like style or safety).
+Finally, we use off-the-shelf LLMs as judges for accessibility and reproducibility; exploring specialized evaluators with stronger reasoning or generative reward models may yield further gains.
 Acknowledgments
@@
 We also thank Robert Vacareanu for valuable early feedback and discussions.
-Appendix A Training dataset breakdown
-In this section, we show the topic distribution breakdown of the medicine and science training dataset and the rubric type distribution breakdown for the corresponding dataset.
-A.1 Medicine training dataset - Overall Data Distribution
+Appendix A Appendix
+A.1 Details of RaR-Medicine dataset
+The following is illustrates an example from the RaR-Medicine dataset with more details about statistics and categorical distribution of rubrics listed in Table 4 and 6.
+RaR-Medicine Example Question.
+A 50-year-old male patient weighs 65 kg with a pH of 7.05, PCO2 of 15 mmHg, HCO3 of 5 mEq/L, and a base deficit of -40 mEq/L.
+How much sodium bicarbonate should be administered in the first 4 hours to correct his metabolic acidosis?
+Reference Answer.
+To address the severe metabolic acidosis in this 50-year-old patient (weight 65 kg) with a pH of 7.05, it is crucial to approach treatment with caution to avoid complications from overcorrection.
+The base deficit is -40 mEq/L, which indicates a significant acid-base disturbance.
+Typically, the formula for bicarbonate administration is: $\text{Bicarbonate required}=\text{Base Deficit}(\text{mEq/L})\times\text{Body Weight}(\text{kg})\times 0.3$ Applying the values: $40\,\text{mEq/L}\times 65\,\text{kg}\times 0.3=780\,\text{mEq}$ However, it is standard practice to administer a partial correction initially, aiming for safety and gradual improvement.
+Therefore, administering about 150 mEq of sodium bicarbonate over the first 4 hours is a prudent approach.
+This amount aligns with safe clinical guidelines for initial treatment in severe metabolic acidosis, ensuring the patient starts on the path to stabilization without risking harm from rapid correction.
+Rubric (checklist and weights).
+Bicarbonate Calculation (w=5): Essential Criteria: The response must correctly identify and apply the formula (Base Deficit × Body Weight × 0.3) to determine the bicarbonate requirement.
+Safe Dosing Recommendation (w=5): Essential Criteria: The response must state a clear recommendation of administering about 150 mEq of sodium bicarbonate over the first 4 hours.
+Partial Correction Justification (w=4): Important Criteria: The response should explain that only a partial correction is administered initially to avoid complications from rapid overcorrection.
+Step-by-Step Calculation (w=3): Important Criteria: The response must detail the calculation steps, showing that 40 mEq/L × 65 kg × 0.3 equals 780 mEq, before noting the adjusted dose for safety.
+Base Deficit Interpretation (w=2): Optional Criteria: The response may mention that the base deficit of -40 mEq/L indicates severe metabolic acidosis requiring cautious treatment.
+Patient Data Accuracy (w=3): Important Criteria: The response must accurately incorporate the patient’s weight of 65 kg and his critical pH, PCO2, and HCO3 values into the explanation.
+Avoid Overcorrection Risk (w= -1): Pitfall Criteria: Does not mention the risks associated with rapid overcorrection of metabolic acidosis if only the full calculated bicarbonate amount is administered.
 Metric Value
-Total examples 20 166
+Total examples 20,166
 Avg. rubrics per question 7.5
 Avg. question length (words) 45.0
-Table 5: Aggregate statistics for the full medical (train and validation) dataset.
+Table 4: Aggregate statistics for the RaR-Medicine dataset (train and validation) dataset.
 Rubric Type Count Percent
-Important 52 748 34.1
-Essential 47 584 30.7
-Optional 34 261 22.1
-Pitfall 20 215 13.1
+Important 52,748 34.1
+Essential 47,584 30.7
+Optional 34,261 22.1
+Pitfall 20,215 13.1
 Table 5: Rubric-type distribution across all 20,166 examples.
 Topics Count Percent
-Total examples 20 166 100.0
-Medical Diagnosis 10 147 50.3
-Medical Treatment 3 235 16.0
-Medical Knowledge 2 557 12.7
-Medical Diag. and Mngmnt 2 033 10.1
+Total examples 20,166 100.0
+Medical Diagnosis 10,147 50.3
+Medical Treatment 3,235 16.0
+Medical Knowledge 2,557 12.7
+Medical Diag. and Mngmnt 2,033 10.1
 Medical Biology 770 3.8
@@
 Table 6: Distribution of topics in the medical training and validation dataset
-A.2 STEM Dataset - Overall Data Distribution
+A.2 Details of RaR-Science dataset
+This section shows an illustrative example from the RaR-Science dataset with more details about statistics and categorical distibution of rubrics listed in Table 7, 8, 9.
+RaR-Science Example Question.
+Determine the solubility of boric acid ($\mathrm{H_{3}BO_{3}}$) in ethanol ($\mathrm{C_{2}H_{5}OH}$) compared to its solubility in benzene ($\mathrm{C_{6}H_{6}}$), considering the principles of ’likes dissolve likes’ and the role of $K_{sp}$ values.
+Explain your reasoning and provide examples of how immulcifiers could affect the solubility of substances in different solvents.
+Reference Answer.
+Boric acid is more soluble in ethanol than in benzene.
+Rubric (checklist and weights).
+Correct Solubility Direction (w=5): Essential Criteria: The response must clearly identify that boric acid is more soluble in ethanol than in benzene.
+Polarity Principle (w=5): Essential Criteria: The answer should explain how the ’like dissolves like’ principle applies by contrasting the polar nature of ethanol with the non-polar character of benzene.
+Ksp Context (w=4): Important Criteria: The response should account for the role of Ksp values, discussing their typical relevance to solubility even though boric acid is a covalent compound rather than an ionic one.
+Immulcifier Explanation (w=4): Important Criteria: The answer should explain how immulcifiers could modify solubility, providing an example of their effect on solvation in different solvents.
+Chemical Properties (w=4): Important Criteria: The response should analyze the inherent chemical properties of boric acid and the solvents to justify the observed solubility differences.
+Avoid Ionic Assumptions (w= -1): Pitfall Criteria: The answer must not incorrectly assume that Ksp values for ionic compounds directly determine the solubility of a covalent acid such as boric acid.
+Enhanced Detail (w=2): Optional Criteria: The response may include additional examples or a concise explanation of solvation dynamics to further illustrate how solubility is influenced.
 Metric Value
-Total examples 20 625
+Total examples 20,625
 Avg. rubrics per question 7.5
@@
 Rubric Type Count Percent
-Important 52 315 34.8
-Essential 42 739 28.4
-Optional 33 622 22.3
-Pitfall 21 808 14.5
+Important 52,315 34.8
+Essential 42,739 28.4
+Optional 33,622 22.3
+Pitfall 21,808 14.5
 Table 8: Rubric‐type distribution across all 20 625 examples.
+A.3 Training Details
+The training hyperparameters are described in Table 10.
 Topics Count Percent
@@
 Table 9: Distribution of topics in the STEM training and validation dataset
-Appendix B Synthetic Preference Set Generation
-We leverage the publicly-released HealthBench Arora et al.
-(2025) corpus, which contains 5,000 health-related prompts accompanied by expert-written answers.
-Of these, 4,203 datapoints already include an ideal completion vetted by licensed clinicians.
-For every such prompt–ideal pair we automatically generate a perturbed counterpart using o3 with the structured template shown below.
-The template forces the model to (i) spell out a [reasoning] plan for degrading quality, (ii) emit the degraded [perturbed_completion], and (iii) log exact [chunks_added] and [chunks_removed].
-Perturbations are accepted only after manual screening confirms that they are objectively worse, along at least one axis of medical accuracy, completeness, clarity, safety, specificity, structure, or tone, while remaining coherent and free of dangerous advice.
-This procedure produces a balanced evaluation set of 4,203 preferred and 4,203 perturbed responses (8,406 total), which we use in the rubric-versus-Likert experiments in Section 5.
-Appendix C Judge Quality impacts on Post-training
-We assess whether rubric-guided evaluation improves judge effectiveness compared to rubric-free Likert scoring when used for GRPO training.
-Table 11 reports judge accuracy on synthetic medical data, with all policies trained using Qwen2.5-7B and varying judge models.
-Judge Model RaR-Implicit Simple-Likert
-GPT-4o-mini 0.279 0.253
-Qwen-32B-Instruct 0.262 0.254
-Qwen-14B-Instruct 0.250 0.249
-Qwen-7B-Instruct 0.267 0.220
-Table 11: Judge quality comparison: rubric-based evaluation vs pure Likert scoring on synthetic medical rubrics.
-Rubric-based evaluation consistently yields stronger policies across all judge sizes.
-The most pronounced improvement appears with Qwen-7B-Instruct (+0.047), where rubric guidance lifts it from weakest to nearly matching larger models.
-Additionally, rubric-based scores are more tightly clustered (0.250–0.279) than those from Likert-only judges (0.220–0.254), indicating improved consistency.
-These results suggest that rubrics help smaller judges approximate high-quality supervision by breaking evaluation into interpretable, binary criteria.
-This structured approach mitigates scale-related limitations, enabling more reliable reward modeling even with limited-capacity evaluators.
-Appendix D Training Details
-The training hyperparameters are described in Table 12.
 Hyperparameters
@@
 Table 10: GRPO hyperparameter settings for Medical and Science domains.
-Appendix E LLM-Judge Prompts
-Prompt for RUBRIC-LIKERT-IMPLICIT Method System Prompt: You are an expert evaluator.
+Judge Model RaR-Implicit Direct-Likert
+GPT-4o-mini 27.9% 25.3%
+Qwen-32B-Instruct 26.2% 25.4%
+Qwen-14B-Instruct 25.0% 24.9%
+Qwen-7B-Instruct 26.7% 22.0%
+Table 11: Judge quality comparison: rubric-based evaluation vs pure Likert scoring on synthetic medical rubrics.
+Rubric-based evaluation consistently yields stronger policies across all judge sizes.
+The most pronounced improvement appears with Qwen-7B-Instruct (+0.047), where rubric guidance lifts it from weakest to nearly matching larger models.
+Additionally, rubric-based scores are more tightly clustered (0.250–0.279) than those from Likert-only judges (0.220–0.254), indicating improved consistency.
+These results suggest that rubrics help smaller judges approximate high-quality supervision by breaking evaluation into interpretable, binary criteria.
+This structured approach mitigates scale-related limitations, enabling more reliable reward modeling even with limited-capacity evaluators.
+A.4 Evaluation Prompts
+GPQA Evaluation Prompt Determine whether the following model response matches the ground truth answer.
+## Ground truth answer##: Option {correct_answer} or {correct_answer_text}
+## Model Response ##: {response_text}
+A response is considered correct if it’s final answer is the correct option letter (A, B, C, or D), or has the correct answer text.
+Please respond with only "Yes" or "No" (without quotes). Do not include a rationale.
+A.5 Predefined Static Rubrics
+Predefined Static Rubrics for RaR-Static Method The response contains correct information without factual errors, inaccuracies, or hallucinations that could mislead the user.
+The response fully answers all essential parts of the question and provides sufficient detail where needed.
+The response is concise and to the point, avoiding unnecessary verbosity or repetition.
+The response effectively meets the user’s practical needs, provides actionable information, and is genuinely helpful for their situation.
+A.6 LLM-Judge Prompts
+Prompt for RAR-IMPLICIT Method System Prompt: You are an expert evaluator.
 Given a user prompt, a generated response, and a list of quality rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics.
@@
 The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10.
-Example response: ‘‘‘json
+Example response:
+‘‘‘json
 "rating": 7
 User Prompt Template: Given the following prompt, response, and rubrics, please rate the overall quality of the response on a scale of 1 to 10 based on how well it satisfies the rubrics.
-<prompt> prompt </prompt> <response> response </response> <rubrics> rubric_list_string </rubrics> Your JSON Evaluation:
-Prompt for SIMPLE-LIKERT Baseline System Prompt: You are an expert evaluator.
+<prompt>
+{prompt}
+</prompt>
+<response>
+{response}
+</response>
+<rubrics>
+{rubric_list_string}
+</rubrics>
+Your JSON Evaluation:
+Prompt for DIRECT-LIKERT Baseline System Prompt: You are an expert evaluator.
 Given a user prompt and a generated response, please rate the overall quality of the response on a scale of 1 to 10, where 1 is very poor and 10 is excellent.
@@
 The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10.
-Example response: ‘‘‘json
+Example response:
+‘‘‘json
 "rating": 8
 User Prompt Template: Given the following prompt, and response, please rate the overall quality of the response on a scale of 1 to 10.
-<prompt> prompt </prompt> <response> response </response> Your JSON Evaluation:
+<prompt>
+{prompt}
+</prompt>
+<response>
+{response}
+</response>
+Your JSON Evaluation:
 Prompt for REFERENCE-LIKERT Baseline System Prompt: You are an expert evaluator.
@@
 The JSON object should contain a single key "rating" and the value should be an integer between 1 and 10.
-Example response: ‘‘‘json { "rating": 8 }‘‘‘ User Prompt Template: Given the following prompt, reference response, and generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference.
-<prompt> prompt </prompt> <reference_response> reference </reference_response> <generated_response> response </generated_response> Your JSON Evaluation:
-Appendix F Evaluation Prompts
-GPQA Evaluation Prompt Determine whether the following model response matches the ground truth answer.
-## Ground truth answer## Option correct_answer or correct_answer_text ## Model Response ##: response_text A response is considered correct if it’s final answer is the correct option letter (A, B, C, or D), or has the correct answer text.
-Please respond with only "Yes" or "No" (without quotes).
-Do not include a rationale.
-Appendix G Predefined Rubrics
-Predefined Rubrics for RaR-Predefined Method The response contains correct information without factual errors, inaccuracies, or hallucinations that could mislead the user.
-The response fully answers all essential parts of the question and provides sufficient detail where needed.
-The response is concise and to the point, avoiding unnecessary verbosity or repetition.
-The response effectively meets the user’s practical needs, provides actionable information, and is genuinely helpful for their situation.
-Appendix H Synthetic Rubric Generation Prompts
+Example response: ‘‘‘json
+"rating": 8
+User Prompt Template: Given the following prompt, reference response, and generated response, please rate the overall quality of the generated response on a scale of 1 to 10 based on how well it compares to the reference.
+<prompt>
+{prompt}
+</prompt>
+<reference_response>
+{reference}
+</reference_response>
+<generated_response>
+{response}
+</generated_response>
+Your JSON Evaluation:
+A.7 Synthetic Preference Set Generation
+We leverage the publicly-released HealthBench [2] corpus, which contains 5,000 health-related prompts accompanied by expert-written answers.
+Of these, 4,203 datapoints already include an ideal completion vetted by licensed clinicians.
+For every such prompt–ideal pair we automatically generate a perturbed counterpart using o3 with the structured template shown below.
+The template forces the model to (i) spell out a [reasoning] plan for degrading quality, (ii) emit the degraded [perturbed_completion], and (iii) log exact [chunks_added] and [chunks_removed].
+Perturbations are accepted only after manual screening confirms that they are objectively worse, along at least one axis of medical accuracy, completeness, clarity, safety, specificity, structure, or tone, while remaining coherent and free of dangerous advice.
+We further exclude the prompts from HealthBench-1k used for ablations.
+This procedure produces a balanced evaluation set of 3,027 preferred and 3,027 perturbed responses (6,054 total), which we use in the rubric-versus-Likert experiments in Section 5.
+The prompt used for this generation is detailed in Figure A.9.
+A.8 Judge Quality impacts on Post-training
+We assess whether rubric-guided evaluation improves judge effectiveness compared to rubric-free Likert scoring when used for GRPO training.
+Table 11 reports judge accuracy on synthetic medical data, with all policies trained using Qwen2.5-7B and varying judge models.
 Prompt for Synthetic Rubrics Generation: Medical Domain You are an expert rubric writer.
@@
 The reference answer is an ideal response but not necessarily exhaustive; use it only as guidance.
-Appendix I Perturbed Dataset Generation
+A.9 Perturbed Dataset Generation
 Prompt for Health Response Perturbation You are tasked with creating perturbed versions of health-related responses to generate training data for evaluating model robustness.